### v1

# WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada

Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across 13 key themes. Evaluating both Vision Language Models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available, by up to 23%. Our findings highlight the enduring importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
WildFireCan-MMD: A Multimodal dataset for Classification of User-generated Content During Wildfires in Canada
Braeden Sherritt Carleton University                        Isar Nejadgholi National Research Council Canada                        Marzieh Amini Carleton University
1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoringGhanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years. Such data contains crucial details such as reports of casualties, infrastructure damage, urgent needs of affected individuals, and information about missing persons, among other critical information Houston et al. (2015); Alam et al. (2018). Using the shared information effectively can facilitate humanitarian aid efforts for more efficient management and faster recovery. This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop machine learning models trained for sifting through this information.
To foster research in this area, Alam et al. (2018) have developed the CrisisMMD dataset to harness multimodal social media data for disaster response and humanitarian aid. This dataset was collected from Twitter and included earthquakes, hurricanes, wildfires, and floods occurring in 2017. However, the classes of CrisisMMD are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Various information, such as the nature of the damages and the type of assistance required, often differs depending on the specific disaster and its geographic context.
To address this limitation, we focus our work on gathering social media data specifically related to wildfires in Canada. Wildfires are becoming increasingly frequent and severe in this country due to the adverse effects of climate change. Although the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context. Therefore, wildfires in Canada remain underexplored in social media analysis research. Upon inspecting our wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of wildfire-related social media content. To address this, we devised thirteen tailored categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts.
Then, we explored the feasibility of training multimodal classification models for triaging tweets during wildfires. In recent years, the research community has extensively worked on developing multimodal machine learning models for processing and categorizing social media data during natural disasters such as wildfires, hurricanes, and floods Ofli et al. (2020); Dwarakanath et al. (2021); Algiriyage et al. (2022); Basit et al. (2023); Koshy and Elango (2023); Hossain et al. (2022); Rezk et al. (2023); Khattar and Quadri (2022). Almost all of these works use the CrisisMMD dataset for their research, highlighting the scarcity of datasets on this topic.
To Benchmark the available models on WildFireCan-MMD, we evaluate the zero-shot classification capabilities of VLMs. Specifically, we prompt LLaVA (open-source) and gpt-4o-mini (proprietary) for label generation, and benchmark their performance against custom-trained multimodal classifiers using concatenative early and late fusion of RoBERTa and ViT modelsOfli et al. (2020); Basit et al. (2023). While VLMs offer a fast, label-free alternative, our results show that fine-tuned fusion models significantly outperform them. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code is available at WildFireCanMMD-Github.  [^1] Please contact the authors for more information about data sharability.. The contributions of this study are summarized as follows:
  • We developed a 13-category taxonomy through multimodal topic modleing of wildfire-related user-generated content.
  • We introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, labeled based on our taxonomy.
  • Uisng WildFireCan-MMD as benchamrk, we evaluate zero-shot VLMS on this task, finding that while closed-source models outperform open ones, both fall short without fine-tuning, which highlights the need for curated training data. Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
2 Background and Literature Review
In this section, we review the literature on social media analysis for disaster response beyond the machine learning field, as well as the literature on the technical aspects of this project.
2.1 Social media analysis for disaster response
The growing interest among researchers in the role of social media during disasters has led to significant contributions in this field. Reference Fauzi (2023) conducted a bibliometric analysis to explore the use of social media in disaster management. This work provides a detailed review of the field, identifying key publications and future research directions, such as the implications of climate emergencies, integrating big data analytics, and managing knowledge in disaster contexts. Similarly, Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. Furthermore, Algiriyage et al. (2022) offered a systematic review of the application of deep learning in disaster response. They discuss the successes and challenges faced in this area, propose guidelines for future research, and consider the importance of using multimodal data in disaster scenarios. These contributions highlight the significant and varied ways social media can support disaster response efforts. Also, Acikara et al. (2023) explored how social media analytics can improve disaster response by enabling the rapid collection of time-sensitive information from affected areas. They noted the value of social media users as "citizen sensors" for quickly gathering essential data. Their review suggests that using social media analytics in disaster management, particularly for natural and human-made hazards, is a developing area that emphasizes the importance of human-centric insights, such as collective intelligence and location awareness. They also identify a gap in research on managing transport networks in real-time after disasters, which could enhance community resilience and sustainability. In other works, (Dwarakanath et al. (2021), Bukar et al. (2020), Mundottukandi et al. (2024)) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Their reviews offer guidance for selecting research topics and aiding emergency teams in making informed decisions.
Our work is motivated by this literature and focuses on the use of machine learning models to triage social media data to potentially be used in all phases of disaster management.
2.2 Multimodal classification for disaster response
Expanding beyond text-based analysis of social media content, recent studies have emphasized multimodal analysis. We review this literature based on the available datasets and previously explored algorithms.
Datasets: Reference Dong et al. (2021) aimed to improve disaster relief by analyzing social media data to understand public attitudes and demand for relief supplies across different disaster types, using a textual Twitter dataset. They apply both qualitative and quantitative methods, including eight machine learning models, to assess public perception, offering insights for more effective disaster management. Also, Ochoa and Comes (2021) introduced a machine learning model that uses multimodal data, including satellite imagery and social media, to predict housing and shelter needs following disasters, demonstrating the model’s accuracy in forecasting such requirements. The lack of labeled social media imagery data for disaster response has been addressed by Alam et al. (2018), who created CrisisMMD, a comprehensive multimodal dataset from Twitter covering various natural disasters. This dataset, with thousands of manually annotated tweets and images from events like earthquakes, hurricanes, wildfires, and floods, provides valuable annotations for crisis management tasks. The dataset has been extensively utilized in research, shedding light on the complex communication patterns during disasters and the importance of integrating textual and visual data to better understand social media’s role in disaster contexts. The work by Ofli et al. (2020) split the problem into two distinct tasks. ’Task 1’ being informative or non-informative, to determine at a high level whether a given tweet is useful for humanitarian aid purposes. ’Task 2’ being the humanitarian categories or classes which the tweets are sorted into, for which they had a total of eight.
While previous works used a generalized scheme of annotations across various disasters, our work develops a wildfire-specific taxonomy of thirteen classes and introduces, WildFireCan-MMD, a Canadian dataset annotated with these labels.
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the work by Ofli et al. (2020). They found that using a combination of AlBERTLan et al. (2020) as a text embedder, and DeITTouvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Reference Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of the traditional LSTM. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTaLiu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018).
We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMDintroduced in this work.
2.3 Large Vision-Language Models
Recent advances in multimodal LLMs have enabled models to process image and text jointly, expanding their utility across a range of complex tasks. A comprehensive survey by Zhang et al. (2024) outlines the design, training strategies, and benchmark performance of 26 MM-LLMs, while Li et al. (2022) categorizes the evolution of VLMs into task-specific models, vision-language pretraining approaches, and those leveraging large-scale, weakly labeled image-text data. These VLMs now support zero- and few-shot generalization across diverse domains.
In this work, we evaluate gpt-4o-mini and LLaVA as zero-shot classifiers on the WildFireCan-MMD dataset. gpt-4o-mini, OpenAI’s most recent transformer-based VLM at the time Achiam et al. (2023), is closed-source and accessible only via API, limiting transparency around its architecture and training. In contrast, LLaVA Liu et al. (2023) is fully open-source and trained using GPT-4-generated responses paired with user-generated prompts and CLIP-based visual features Radford et al. (2021). It employs visual instruction tuning to align vision and language modalities and has shown strong performance on multimodal instruction-following benchmarks. The release of models like LLaVA reflects broader efforts to build high-performing, publicly accessible VLMs in response to the closed nature of commercial systems like GPT-4.
3 The WildFireCan-MMD Dataset
3.1 Data Collection
We used X’s Pro-Tier paid API to gather a diverse dataset of posts containing information, updates, and discussions related to the British Columbia/Alberta wildfires of 2022 and 2023, and the Jasper wildfires of 2024. This dataset contains tweets between the months of April and June for each year and consists of 4,688 unique image-text pairs. We collected this data using the Tweepy library and conducted a hashtag-based search, using hashtags we found to yield good results through experimentation. This was a crucial part of our data collection as we wanted to ensure we gathered tweets originating from Canada, that were also on topic. Due to not many users in Canada having their specific geo-location turned on for Twitter, we couldn’t simply scrape by region, or very little data would be returned. As a result, we had to adopt a hashtag-based search. Due to the uniqueness of the hashtags, which are only really used in Canada during wildfires, this yielded a dataset of relevant tweets with mostly users from the specified region. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated. Table A.1 presents the full queries used for data collection.
3.2 Developing a Taxonomy
To examine the content of the tweets and discover the hidden, latent semantics of our data, we used the BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al. (2022); Egger and Yu (2022); Medvecki et al. (2023). We then manually inspected the topics to identify the categories. In doing so, we considered the stakeholders who will be using this data. Implementation details of the BERTopic clustering are provided in Appendix B. In the end, we selected 13 classes which were important and recurrent enough to be considered categories for this dataset. The full list of categories and their descriptions is given in Table 1.
The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios. For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel. The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality. The ‘Political’ category could be useful to authorities to gauge public opinion on events surrounding the disaster.
Table 1: Categories of WildFireCan-MMD, their descriptions and the size of each class.
Category Description Class Size
Evacuees Information relating to evacuees, their movements, needs, location. 242
Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1070
General Information General info about the wildfire situation, such as total hectares burned. 186
Preparedness Information for the general public to prepare themselves and their property. 265
Weather Reports Information relating to the weather with a specific location mentioned. 316
Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 620
Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 378
Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 293
Political Posts directed towards political figures or parties (excluding situation updates). 322
Support Mental health and financial support, temporary housing for livestock. 164
Insurance Information relating to insurance, employment insurance, and EI benefits. 142
Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 116
Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 569
3.3 Data Labeling
We labeled the collected dataset according to the taxonomy of categories presented in Table 1. Labelling was done in-house by one of the authors, using Label Studio [^2] https://labelstud.io/. Since this task is not subjective, we did not consider multi-annotator labeling. Label Studio was chosen for its ability to search through the text fields of the dataset for specific keywords, rather than running through the dataset in sequence, which made labeling easier. For example, this allowed us to search for ’air quality’ and see all posts containing these keywords. This allowed us to focus on one category at a time and get the majority of a category labeled before moving on to another. When searching by keywords was no longer possible, labeling in sequence resumed.
Label: Reports of Actions of Responders
(a) Text: PAFD is on the scene of a fire in Roger Creek Park. There is no threat to structures at this time. #portalberni #BCWildfire
Label: Infrastructure
(b) Text: The wildfires in #AB have burned down this schoolhouse. #ABWildfires
Label: Warnings & Status Updates
(c) Text: Wildfire updates, July 24: Fire reaches Jasper townsite | ’Significant loss has occured’: Parks Canada | Firefighters working to protect critical infrastructure #abfire
Figure 1: Examples of labeled posts from WildFireCan-MMD.
Examples of labeled tweets are presented in Figure 1. Figure 2 displays the class distribution in WildFireCan-MMD. Also, for training and testing the custom classifiers, given in Section 4, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets. The size of labeled classes are presented in Table 1.
Figure 2: Class distribution of WildFireCan-MMD.
4 Classification Methodology
This section describes a range of algorithms evaluated for the 13-class classification of WildFireCan-MMD, utilizing text and image modalities. We also experimented with using Optical Character Recognition (OCR) to further extract more information from the images and create another channel of text to aid in classification. We experiment with using VLMs as zero-shot classifiers, and compare these results with fine-tuned transformer-based classifiers as well as to baseline models such as Random Forest, and Gaussian Naive Bayes. We present the implementation details of classification algorithms in the following.
4.1 VLM as Zero-Shot Classifiers
We crafted classification prompts for both gpt-4o-mini and LLaVa in an ’all-at-once’ method, meaning that with each query to the model, a list of all categories and their descriptions are given as a prompt, along with a sample (text + image) from the data to classify. We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter. The system and user prompts can be seen in Table C.1.
4.1.1 gpt-4o-mini:
We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18. A specified content field in each query contains our system prompt, text from the Tweet, and user prompt. The image for the tweet is also sent with this payload.
The ’temperature’ is specified as being $0.1$ for reproducibility. ’max_tokens’ is set to be $1024$, as the text plus the image plus the model’s response will never be greater than this. The responses are saved as a JSONL file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1. This then allowed us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
4.1.2 LLaVA:
As an open source model, querying LLaVA required a different approach than gpt-4o-mini. For our work, we opted for the 13B 4-bit version of LLaVa. The temperature was again set to $0.1$, and max_tokens set to $1024$. A custom conversation template was created using the original LLaVA codebase to construct our system prompt using the roles USER and ASSISTANT. The question file was created using the steps discussed above for gpt-4o-mini, the only difference being the addition of ’<>’ characters around each option as recommended by the source [^4] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md for querying LLaVA with a multiple-choice prompt. The text from the tweet is passed similarly to the method described above for gpt-4o-mini, and the image is converted to a tensor and passed with this to the model. The answers are saved similarly as described for gpt-4o-mini and saved, allowing us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
4.2 Custom Classifiers
To create a custom multimodal classifier for our dataset, the following steps were taken to implement concatenative early and late fusion.
Early and late fusion 2-head structure::This model consists of two architectures working in parallel, an image head and a text head. The image head uses the ViT transformer, and the text head uses the RoBERTa transformer. Another linear layer is added for prediction. Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion. We created another model with the same strategy but utilizing late fusion instead, where each head has its own final linear layer and functions as an independent classifier. The two heads output probabilities, which are then averaged for a final result.
Training and Evaluation: A random seed was used for training reproducibility. Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by google/vit-base-patch16-384), and a max length for the text of 200. Twitter posts can have a maximum length of 200 characters, but after viewing the average and max length of all posts within the dataset, we learned that most users don’t go anywhere near this limit, and so this value could be decreased. We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss. Our experimentation found that the model converged at around 10-15 epochs. The model was then tested on the test set (9̃00 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
Additional Information with OCR: Trying to improve our classification metrics without gathering and labeling more data, we considered other ways in which we could extract more information from the data we had. Something we noticed in much of the data, specifically the Evacuees, Smoke & Air Quality, General Information, Preparedness, Warnings and Status Updates, and Weather Reports classes, was that many of the images contained valuable text which we believed could aid in classification. Although ViT and RoBERTA both use the transformer architecture, the way text is processed differs from images and doesn’t allow for ViT to differentiate text within images from the surrounding pixels in this current implementation.
This led us to consider ways in which we could extract text from the images in our dataset using the EasyOCR tool [^5] https://github.com/JaidedAI/EasyOCR. EasyOCR works by taking an image, preprocessing it, then applying the CRAFT algorithm to find bounding-box regions of text in an image. The pixels in these regions are sent to a text recognition model composed of a CRNN with three components: feature extraction, sequence labeling, and decoding to extract the text from these regions and turn them into strings as output. Deploying this model on our dataset, we extracted all of the text from every image and saved it to a new column as a pre-processing step. As a result, our dataset contains four columns: text, extracted_text, images, and labels. If an image was found to not have any text, a filler sentence was used: ’Image does not contain text.’.
Utilizing this extra column of text data, we added another RoBERTa head to our classifier. This new classifier, now with three separate heads concatenated and fed into a single linear layer, as appeared in Figure D.2. Note that this figure represents the early fusion version. A late fusion version was also created similarly as described for the 2-head classifier. The model was then trained with the same settings as described previously, for a similar number of epochs.
5 Experimental Results
Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and two baseline models per class and as an average over all classes. Also, figure E.1 shows the confusion matrices of the VLMs and the custom classifiers. Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
Table 2: F1-scores per-class for VLMs, custom classifiers, and baseline models.
VLMs Custom Classifiers Baseline Classifiers
Class LLaVa gpt-4o-mini 2-head-early 3-head-early 2-head-late 3-head-late R. Forest Gaussian NB
Evacuees 0.17 0.37 0.77 0.73 0.79 0.74 0.29 0.50
General Information 0.14 0.16 0.59 0.46 0.53 0.59 0.35 0.67
Preparedness 0.61 0.65 0.82 0.81 0.86 0.84 0.52 0.80
Weather Reports 0.59 0.64 0.88 0.91 0.91 0.91 0.45 0.66
Warnings & Status Updates 0.36 0.45 0.80 0.76 0.77 0.74 0.50 0.57
Reports of Actions of Responders 0.29 0.63 0.89 0.86 0.85 0.83 0.48 0.74
Infrastructure 0.38 0.40 0.92 0.95 0.90 0.93 0.62 0.64
Political 0.70 0.77 0.84 0.84 0.84 0.85 0.28 0.78
Insurance 0.93 0.90 0.93 0.89 0.93 0.91 0.92 0.91
Advertisement 0.08 0.68 0.76 0.76 0.74 0.75 0.23 0.81
Smoke and Air Quality 0.76 0.79 0.86 0.86 0.86 0.85 0.58 0.70
Support 0.25 0.55 0.73 0.63 0.75 0.78 0.45 0.63
Other 0.05 0.54 0.78 0.78 0.76 0.79 0.36 0.60
Average 0.41 0.58 0.81 0.79 0.81 0.81 0.46 0.69
Weighted Average 0.45 0.60 0.83 0.81 0.82 0.82 0.48 0.67
LLaVa has limited capabilities.Testing on WildFireCan-MMD, We observed that LLaVa is not able to reach an acceptable result, performing close to the Random Forest baseline and underperforming gpt-4o-mini by a large margin. Specifically, we noted a large difference in F1-score between gpt-4o-mini and LLaVA for the Advertisement and Other classes (gpt4o-mini achieved between 0.5 and 0.6, whereas LLaVA reaches only less than 0.1). This can be explained by the fact that these two classes have the highest variance by far and are hard to handle for LLaVa. Specifically, the Other category is very diverse since tweets are labeled as so if they don’t fit into any of the other categories. Referring to confusion matrices (Figure E.1(a)), we can see that LLaVA struggles with labeling the Warnings & Status Updates class, where it often classified tweets from other classes as such, especially the Other class.
gpt-4o-mini shows strengths and limitations across various categories. On the other hand, gpt-4o-mini, being much larger and more sophisticated than LLaVA, outperforms it in almost every class except for the Insurance category, where it has been beaten by LLaVA by a couple of percent. Interestingly, gpt-4o-mini reaches a similar F1-score as custom classifiers on the Smoke & Air Quality, Advertisement, Insurance and Political classes, which combined encompass over a third of the dataset. This is probably due to the lower variance of these classes. On the contrary, the classes in which it suffers the most are Evacuees, General Information, Warnings & Status Updates, Infrastructure which contain some of the highest variance. It is difficult to construct a prompt, which encompasses the variance of these classes, while staying within the token limit. So, this may be one reason why gpt4-o performed worse than custom models in these categories. As seen in Figure E.1(b) we can see that gpt-4o-mini had the most trouble with the Warnings & Status Updates class, where it often confused it with the General Information and Evacuees classes. This effect caused poor performance in the Infrastructure class as well.
Custom classifiers are more robust and reliable. Referring to Table 2, we can see that the two and three-head classifiers provide similar results. Our 2-head classifiers have an average weighted F1-score of 0.83 for early fusion, and 0.82 for late fusion. The 3-head classifiers have an average weighted F1-score of 0.81 for early fusion, and 0.82 for late fusion. Surprisingly, the three-head classifier performs slightly worse in most classes than the two-head classifier, even though it receives extra contextual information from the images. One important consideration, however, when thinking about the utility and purpose of this work, is that the 3-head classifiers perform slightly better on key classes such as Infrastructure, Weather Reports, Smoke & Air Quality, General Information, and Support. Infrastructure is an important class because it can provide information to first responders and administrative personnel about key infrastructure that has been damaged or is currently in need of focus. Weather Reports is also important because it can provide insights and predictions about weather patterns and likely areas of new fires. General Information is useful to administrators and the general public as this class contains broad information about the situation as a whole.
6 Discussion
This work benchmarks VLMs on a new dataset. A critical consideration in evaluation of the VLMs is that WildFireCan-MMD establishes a novel benchmark for evaluating these models. Most evaluation research works are performed with pre-existing datasets, and these datasets are also highly likely to be used for training VLMs and do not represent a genuine test set. Our newly created dataset evaluates the real generalization capabilities of such models since it was not available to the creators of these models during training. This is important for information management during crises because new disasters have unique requirements.
The use of VLMs involves tradeoffs. Practical implications are involved in benchmarking with VLMs in a zero-shot setting. We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles. The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment. Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al. (2024). On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management. If performance is improved, the open-source nature of this model might make it suitable for disaster response, allowing compliance with responsible AI guidelines through offering access to the code base and ensuring ethical scrutiny. Particularly, the transparent nature of open source is crucial when handling sensitive data that may include personal information. Despite gpt-4o-mini performing better on our test set, it is important to keep these above considerations in mind. When labeled data is unavailable, VLMs can function as a zero-shot classifier, but in practice, neither model is particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
Fine-tuning on task-specific data is essential. Results presented in Table 2 demonstrate that even simple approaches such as Random Forest and Gaussian NB, when trained with task-specific labelled data, provide similar or better results than what is achieved from zero-shot VLMs. Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier. Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning advanced classification structures, such as pre-trained transformer-based text and image encoders.
The custom classifiers we have developed serve as a foundational step for triaging tweets during wildfires, providing an initial performance benchmark and demonstrating our approach’s feasibility. However, to enhance the accuracy and robustness of our model, future work will need to adopt a more comprehensive approach to experimentation. One area for improvement is the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
7 Conclusion
General-purpose VLMs, while powerful, struggle to perform reliably in high-stakes, context-rich scenarios like disaster response. To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts (text + image) curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation. Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs. As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
Limitations
While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
The fisrt limitation is the scope and size of the dataset. Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate general-purpose models. This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
Platform and geographic bias might be present in our data. The dataset is sourced exclusively from the X platform and focuses on Canadian wildfires. This introduces both platform-specific and regional biases in language use, imagery, and public discourse. As a result, models trained on this dataset may not perform as well when applied to other social platforms or international wildfire contexts.
Moreover, zero-shot evaluation of VLMs introduces limitations. Prompt engineering for zero-shot setups is a heuristic process, and better prompts might yield different results.
Appendix A Data collection Queries
Table A.1 shows the full queries used to collect data from X.
Year Query
2022/2023 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
Table A.1: The full queries used for scraping data from X.
Appendix B Topic modeling parameters
In order to optimize the performance of BERTopic on our dataset, careful consideration of hyperparameters is key McInnes et al. (2018). First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^6] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model. Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality. We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15. After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors. HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers. HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
After forming the clusters, BERTopic extracts keywords from each topic to uncover their central theme. For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting a set number of topics to sort the data into, as at this stage, we were unsure of what sort of topics were contained in our data. BERTopic had sorted our data into over 100 fine-grained topics. This was still relatively large to manage, and it was difficult to determine a humanitarian category for each fine-grained topic. After some experimentation, we systematically decreased the total number of topics to determine which topics remained over time. Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics. Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders. We also noticed some less useful posts, such as advertisements and political debates/complaints. Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
Using BERTopic in the way described proved to be useful, as it gave us a good understanding from the beginning of what our future categories were to look like. Relying on this method alone was not enough, however, and having to look through the data manually at some point was inevitable. Many rounds of manual annotation ensued on a small portion of the data, after having it randomly mixed, in order to avoid constructing our categories on themes that may only exist at a certain time during the wildfire event. Through performing these steps, keeping in mind our discovered topics through BERTopic and humanitarian goals, new categories were added or dropped based on what was seen in the data. A final round of labelling then commenced, and the entire dataset was manually labeled.
Keywords: [crews, safe, poles, damaged, continue, lines, working restore, power, conditions, restore]
(a) Useful
Keywords: [highway, closed, closures, abroads, detour, road, route, bchwy4, cameron, cameron lake]
(b) Useful
Keywords: [’yycnow’, ’yycbusiness’, ’yycbusiness yycnow’, ’yyc yycbusiness’, ’abfires abwildfire’, ’calgary’, ’startups’, ’wordify’, ’smallbusiness’, ’smallbusiness startups’]
(c) Non-useful
Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
Appendix C Prompts for zero-shot classification
We show the full prompt used to evaluate VLMs in Table C.1.
Table C.1: System and User prompt used for gpt-4o-mini and LLaVa.
System Prompt:
"You are an assistant who is being given an image and text pair as a Twitter post which was created during a natural disaster event. Your task is to use information from both the text and the image to decide which option the post should be labeled as. You must pay close attention to each option when deciding which label to use."
User Prompt:
"Which option should this post be labeled as?
A. Evacuees (information relating to evacuees, their movements, needs, location, etc)
B. General Information (GENERAL facts about the wildfire situation, hectares burned)
C. Preparedness (information for the general public to prepare themselves and property for wildfires)
D. Weather Reports (information relating to the weather, satellite imagery, radar imagery)
E. Warnings & Status Updates (warnings/updates to the public from authoritative bodies, fire bans, specific information relating to a certain time or area)
F. Reports of Actions of Responders (prescribed burns, responders responding to a specific location)
G. Infrastructure (road closures, damaged buildings or property, traffic)
H. Political (mentions of political or public figures or parties)
I. Insurance (mentions of insurance)
J. Advertisement (information about restaurants, food, apps or services)
K. Smoke & Air Quality (information about smoke or air quality, masks, breathing)
L. Support (information about financial, mental health, or other types of support for people)
M. Other (the post does not fit well in one of the previous categories)
You may only answer with the chosen option’s letter."
Appendix D Structures of custom classifiers
Figures D.1 and D.2 show the structures of the finetuned models. We described these models in detail in Section 4.2.
Figure D.1: Custom multimodal 2-head early fusion classifier structure.
Figure D.2: Custom multimodal 3-head early fusion classifier structure.
Appendix E Confusion matrix for all classifiers
For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figure E.1.
a) LLaVA, b) gpt-4o-mini, c) 2head-early, d) 3head-early, e) 2head-late, f) 3head-late
Figure E.1: Confusion matrices of classifiers and VLMs.

### v2

@@
 Social media offers real-time updates, but extracting relevant insights remains a challenge.
-We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across 13 key themes.
-Evaluating both Vision Language Models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available, by up to 23%.
+We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes.
+Evaluating both vision-language models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available.
+Our best-performing transformer-based fine-tuned model reaches 83% f-score, outperforming gpt4 by 23%.
+Also, we demonstrate how this model can be used to uncover trends during wildfires.
 Our findings highlight the enduring importance of tailored datasets and task-specific training.
@@
 (2021); Kumar et al.
-(2024a), and sentiment and behavioral trend analysis Ma et al.
+(2024a), pandemic discourse analysis Imran et al.
+(2022) and sentiment and behavioral trend analysis Ma et al.
 (2024); Curto et al.
 (2024).
-Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years.
-Such data contains crucial details such as reports of casualties, infrastructure damage, urgent needs of affected individuals, and information about missing persons, among other critical information Houston et al.
+Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam et al.
+(2018); Ofli et al.
 (2020); Alam et al.
-(2018).
-Using the shared information effectively can facilitate humanitarian aid efforts for more efficient management and faster recovery.
-This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information.
-Therefore, there is a critical need to develop machine learning models trained for sifting through this information.
-To foster research in this area, Alam et al.
-(2018) have developed the CrisisMMD dataset to harness multimodal social media data for disaster response and humanitarian aid.
-This dataset was collected from Twitter and included earthquakes, hurricanes, wildfires, and floods occurring in 2017.
-However, the classes of CrisisMMD are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios.
-As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked.
-Various information, such as the nature of the damages and the type of assistance required, often differs depending on the specific disaster and its geographic context.
-To address this limitation, we focus our work on gathering social media data specifically related to wildfires in Canada.
-Wildfires are becoming increasingly frequent and severe in this country due to the adverse effects of climate change.
-Although the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context.
-Therefore, wildfires in Canada remain underexplored in social media analysis research.
-Upon inspecting our wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of wildfire-related social media content.
-To address this, we devised thirteen tailored categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts.
-Then, we explored the feasibility of training multimodal classification models for triaging tweets during wildfires.
-In recent years, the research community has extensively worked on developing multimodal machine learning models for processing and categorizing social media data during natural disasters such as wildfires, hurricanes, and floods Ofli et al.
 (2021); Dwarakanath et al.
 (2021); Algiriyage et al.
+(2022); Hossain et al.
+(2022); Khattar and Quadri (2022); Basit et al.
+(2023); Koshy and Elango (2023); Rezk et al.
+(2023).
+Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information.
+The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston et al.
+(2015).
+This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information.
+Therefore, there is a critical need to develop an automated method for sifting through this information.
+To foster research in this area, Alam et al.
+(2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes.
+However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios.
+As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked.
+Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context.
+The HumAID dataset by Alam et al.
+(2021) also contains wildfire data from Canada, but only a couple of thousand samples from 2016.
+Therefore, wildfires in Canada remain underexplored in social media analysis research.
+To address this limitation, we focus our work on gathering social media data specifically related to Canadian wildfires, as they are becoming increasingly frequent and severe due to the adverse effects of climate change.
+Upon initial inspection of our Canadian wildfire-specific social media dataset, we observed that existing categorization schemes do not capture the specificity of this data.
+To address this, we devised a novel taxonomy of twelve categories reflecting the unique information needs and communication patterns of Canadian wildfire contexts.
+Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
+To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several VLMs, as well as the feasibility of training custom classifiers.
+Our best-performing classifier is a fine-tuned model, which uses RoBERTa Liu (2019) and Vision Transformer (ViT) models Ofli et al.
 (2020); Basit et al.
-(2023); Koshy and Elango (2023); Hossain et al.
-(2022); Rezk et al.
-(2023); Khattar and Quadri (2022).
-Almost all of these works use the CrisisMMD dataset for their research, highlighting the scarcity of datasets on this topic.
-To Benchmark the available models on WildFireCan-MMD, we evaluate the zero-shot classification capabilities of VLMs.
-Specifically, we prompt LLaVA (open-source) and gpt-4o-mini (proprietary) for label generation, and benchmark their performance against custom-trained multimodal classifiers using concatenative early and late fusion of RoBERTa and ViT modelsOfli et al.
-(2020); Basit et al.
+(2023) to jointly process text and image parts of tweets.
+While WildFireCan-MMD is sourced from data posted in 2022-2024, we collected an additional set of unlabeled tweets across seven years (2018-2025).
+We then used our best-performing classifier to analyze this data and uncover trends and patterns of social media use in Canada.
+The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at WildFireCanMMD-Github [^1] Please contact the authors for more information about data sharability., under CC-BY-NC license.
+The contributions of this study are summarized as follows:
+  • We devised a twelve-category taxonomy through multimodal topic modeling of Canadian wildfire-related user-generated content, and introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, manually labeled based on our taxonomy.
+  • We compared custom multimodal transformer models trained on WildFireCan-MMD with several VLMs and classical methods.
+Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
+  • We collected an additional 46,279 posts from X, spanning across seven years, used our trained classifier to label them, and reported the uncovered trends and patterns of social media use during wildfires in Canada.
+2 Background and Literature Review
+2.1 Social media analysis for disaster response
+Significant research has been conducted on the role of social media during disasters.
+Fauzi (2023) conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions.
+Acikara et al.
+(2023) highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information.
+Muniz-Rodriguez et al.
+(2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster.
+In other works, Dwarakanath et al.
+(2021), Bukar et al.
+(2020), and Mundottukandi et al.
+(2024) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities.
+Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage et al.
+(2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
+Our work is motivated by this literature and focuses on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
+2.2 Multimodal classification for disaster response
+We review this literature based on the available datasets and previously explored algorithms.
+Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters Olteanu et al.
+(2014); Dong et al.
+(2021); Ochoa and Comes (2021), it is now generally considered that multimodal approaches offer stronger results Ofli et al.
+(2020); Abavisani et al.
+(2020); Ramirez et al.
 (2025).
-While VLMs offer a fast, label-free alternative, our results show that fine-tuned fusion models significantly outperform them.
-The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code is available at WildFireCanMMD-Github.
-[^1] Please contact the authors for more information about data sharability..
-The contributions of this study are summarized as follows:
-  • We developed a 13-category taxonomy through multimodal topic modleing of wildfire-related user-generated content.
-  • We introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, labeled based on our taxonomy.
-  • Uisng WildFireCan-MMD as benchamrk, we evaluate zero-shot VLMS on this task, finding that while closed-source models outperform open ones, both fall short without fine-tuning, which highlights the need for curated training data.
-Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
-2 Background and Literature Review
-In this section, we review the literature on social media analysis for disaster response beyond the machine learning field, as well as the literature on the technical aspects of this project.
-2.1 Social media analysis for disaster response
-The growing interest among researchers in the role of social media during disasters has led to significant contributions in this field.
-Reference Fauzi (2023) conducted a bibliometric analysis to explore the use of social media in disaster management.
-This work provides a detailed review of the field, identifying key publications and future research directions, such as the implications of climate emergencies, integrating big data analytics, and managing knowledge in disaster contexts.
-Similarly, Muniz-Rodriguez et al.
-(2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster.
-Furthermore, Algiriyage et al.
-(2022) offered a systematic review of the application of deep learning in disaster response.
-They discuss the successes and challenges faced in this area, propose guidelines for future research, and consider the importance of using multimodal data in disaster scenarios.
-These contributions highlight the significant and varied ways social media can support disaster response efforts.
-Also, Acikara et al.
-(2023) explored how social media analytics can improve disaster response by enabling the rapid collection of time-sensitive information from affected areas.
-They noted the value of social media users as "citizen sensors" for quickly gathering essential data.
-Their review suggests that using social media analytics in disaster management, particularly for natural and human-made hazards, is a developing area that emphasizes the importance of human-centric insights, such as collective intelligence and location awareness.
-They also identify a gap in research on managing transport networks in real-time after disasters, which could enhance community resilience and sustainability.
-In other works, (Dwarakanath et al.
-(2021), Bukar et al.
-(2020), Mundottukandi et al.
-(2024)) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities.
-Their reviews offer guidance for selecting research topics and aiding emergency teams in making informed decisions.
-Our work is motivated by this literature and focuses on the use of machine learning models to triage social media data to potentially be used in all phases of disaster management.
-2.2 Multimodal classification for disaster response
-Expanding beyond text-based analysis of social media content, recent studies have emphasized multimodal analysis.
-We review this literature based on the available datasets and previously explored algorithms.
-Datasets: Reference Dong et al.
-(2021) aimed to improve disaster relief by analyzing social media data to understand public attitudes and demand for relief supplies across different disaster types, using a textual Twitter dataset.
-They apply both qualitative and quantitative methods, including eight machine learning models, to assess public perception, offering insights for more effective disaster management.
-Also, Ochoa and Comes (2021) introduced a machine learning model that uses multimodal data, including satellite imagery and social media, to predict housing and shelter needs following disasters, demonstrating the model’s accuracy in forecasting such requirements.
-The lack of labeled social media imagery data for disaster response has been addressed by Alam et al.
-(2018), who created CrisisMMD, a comprehensive multimodal dataset from Twitter covering various natural disasters.
-This dataset, with thousands of manually annotated tweets and images from events like earthquakes, hurricanes, wildfires, and floods, provides valuable annotations for crisis management tasks.
-The dataset has been extensively utilized in research, shedding light on the complex communication patterns during disasters and the importance of integrating textual and visual data to better understand social media’s role in disaster contexts.
-The work by Ofli et al.
-(2020) split the problem into two distinct tasks.
-’Task 1’ being informative or non-informative, to determine at a high level whether a given tweet is useful for humanitarian aid purposes.
-’Task 2’ being the humanitarian categories or classes which the tweets are sorted into, for which they had a total of eight.
-While previous works used a generalized scheme of annotations across various disasters, our work develops a wildfire-specific taxonomy of thirteen classes and introduces, WildFireCan-MMD, a Canadian dataset annotated with these labels.
+Alam et al.
+(2018) addressed the shortage of multimodal labeled social media data for disaster response by introducing CrisisMMD, a widely used dataset containing annotated tweets and images from various natural disasters.
+Later extended by datasets like HumAID Alam et al.
+(2021), CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images.
+It remains the most prominent dataset in this domain, used in many recent studies Basit et al.
+(2023); Koshy and Elango (2023); Shetty et al.
+(2024).
+While previous works such as CrisisMMD used a generalized scheme of annotations across various types of different disasters, our work develops a wildfire-specific taxonomy of twelve categories and introduces WildFireCan-MMD, a new Canadian wildfire dataset.
 Algorithms: Previous work by Ofli et al.
@@
 (2023), expands on Ofli et al.
-(2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the work by Ofli et al.
-(2020).
+(2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work.
 They found that using a combination of AlBERT Lan et al.
@@
 (2021) as an image embedder, provided the best results for the eight-class classification task.
-Reference Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of the traditional LSTM.
+Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs.
 Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods.
@@
 (2018).
-We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMDintroduced in this work.
-2.3 Large Vision-Language Models
-Recent advances in multimodal LLMs have enabled models to process image and text jointly, expanding their utility across a range of complex tasks.
-A comprehensive survey by Zhang et al.
-(2024) outlines the design, training strategies, and benchmark performance of 26 MM-LLMs, while Li et al.
-(2022) categorizes the evolution of VLMs into task-specific models, vision-language pretraining approaches, and those leveraging large-scale, weakly labeled image-text data.
-These VLMs now support zero- and few-shot generalization across diverse domains.
-In this work, we evaluate gpt-4o-mini and LLaVA as zero-shot classifiers on the WildFireCan-MMD dataset.
-gpt-4o-mini, OpenAI’s most recent transformer-based VLM at the time Achiam et al.
-(2023), is closed-source and accessible only via API, limiting transparency around its architecture and training.
-In contrast, LLaVA Liu et al.
-(2023) is fully open-source and trained using GPT-4-generated responses paired with user-generated prompts and CLIP-based visual features Radford et al.
+We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
+In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al.
+(2024); Li et al.
 (2022).
-It employs visual instruction tuning to align vision and language modalities and has shown strong performance on multimodal instruction-following benchmarks.
-The release of models like LLaVA reflects broader efforts to build high-performing, publicly accessible VLMs in response to the closed nature of commercial systems like GPT-4.
+Specifically, we evaluate three VLMs, GPT-4 Achiam et al.
+(2023), LLaVA Liu et al.
+(2023), and SmolVLM Marafioti et al.
+(2025), to assess performance across a range of capabilities and computational requirements.
+GPT-4 serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
 3 The WildFireCan-MMD Dataset
 3.1 Data Collection
-We used X’s Pro-Tier paid API to gather a diverse dataset of posts containing information, updates, and discussions related to the British Columbia/Alberta wildfires of 2022 and 2023, and the Jasper wildfires of 2024.
-This dataset contains tweets between the months of April and June for each year and consists of 4,688 unique image-text pairs.
-We collected this data using the Tweepy library and conducted a hashtag-based search, using hashtags we found to yield good results through experimentation.
-This was a crucial part of our data collection as we wanted to ensure we gathered tweets originating from Canada, that were also on topic.
-Due to not many users in Canada having their specific geo-location turned on for Twitter, we couldn’t simply scrape by region, or very little data would be returned.
-As a result, we had to adopt a hashtag-based search.
-Due to the uniqueness of the hashtags, which are only really used in Canada during wildfires, this yielded a dataset of relevant tweets with mostly users from the specified region.
+To construct our dataset, we utilized X’s Pro-Tier paid API to collect a diverse set of posts surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024.
+The dataset comprises 4,688 unique image-text pairs gathered from tweets posted between April and June of each respective year.
+Data collection was conducted using a hashtag-based search strategy.
+Table A.1 presents the full queries used for data collection.
+To curate the hashtags, we started with seed hashtags and used an iterative process of adding new hashtags by collecting sample tweets.
+This approach was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location.
+Instead, we identified and utilized hashtags that, through iterative experimentation, were found to be effective indicators of relevant content originating from Canada during wildfire events.
+The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions.
 The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated.
-Table A.1 presents the full queries used for data collection.
-3.2 Developing a Taxonomy
-To examine the content of the tweets and discover the hidden, latent semantics of our data, we used the BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al.
+3.2 Taxonomy Development
+To examine the content of the tweets and discover the hidden latent semantics of our data, we used BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al.
 (2022); Egger and Yu (2022); Medvecki et al.
 (2023).
-We then manually inspected the topics to identify the categories.
-In doing so, we considered the stakeholders who will be using this data.
-Implementation details of the BERTopic clustering are provided in Appendix B.
-In the end, we selected 13 classes which were important and recurrent enough to be considered categories for this dataset.
+Further explanation of the BERTopic clustering method are provided in Appendix B.
+Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting number of topics to sort the data into.
+When the number of clusters were unspecified, BERTopic had created over 100 fine-grained topics.
+This was relatively large to manage, and difficult to determine a humanitarian category for each fine-grained topic.
+We gradually decreased the number of topics, and for each number, we manually inspected the representative visual samples of topics and keywords generated by BERTopic.
+Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics.
+Some examples of useful topics were posts about infrastructure, weather reports, and reports of actions of first responders.
+We also noticed some less useful posts, such as advertisements and political debates/complaints.
+Some examples of these topics, including representative images and corresponding keywords, generated by BERTopic, can be seen in Figure B.1.
+We then manually inspected the topics and selected 12 classes that were informative and consistently appeared across a varying maximum number of clusters to be considered designated categories for this dataset.
 The full list of categories and their descriptions is given in Table 1.
+We also include an “Other” to accommodate the examples that do not fit into any of these 12 categories.
 The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios.
@@
 The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality.
-The ‘Political’ category could be useful to authorities to gauge public opinion on events surrounding the disaster.
-Table 1: Categories of WildFireCan-MMD, their descriptions and the size of each class.
-Category Description Class Size
+The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
+Table 1: Categories of WildFireCan-MMD, their descriptions and total count of each class.
+Category Description Count
 Evacuees Information relating to evacuees, their movements, needs, location. 242
@@
 We labeled the collected dataset according to the taxonomy of categories presented in Table 1.
-Labelling was done in-house by one of the authors, using Label Studio [^2] https://labelstud.io/.
-Since this task is not subjective, we did not consider multi-annotator labeling.
-Label Studio was chosen for its ability to search through the text fields of the dataset for specific keywords, rather than running through the dataset in sequence, which made labeling easier.
-For example, this allowed us to search for ’air quality’ and see all posts containing these keywords.
-This allowed us to focus on one category at a time and get the majority of a category labeled before moving on to another.
-When searching by keywords was no longer possible, labeling in sequence resumed.
+Over the course of labeling, we ensured that the data did not contain personal information.
+Labeling was carried out in-house by one of the authors using Label Studio [^2] https://labelstud.io/.
+To ensure the reliability of the labeling process, we first discussed the definitions and boundaries of each category among the research team and reached a shared understanding.
+Through this discussion, we determined that the task was not largely subjective, as the categories are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image).
+Therefore, we did not consider multi-annotator labeling necessary.
+Examples of labeled tweets are presented in Figure 1.
+The size of labeled classes are presented in Table 1.
 Label: Reports of Actions of Responders
@@
 Figure 1: Examples of labeled posts from WildFireCan-MMD.
-Examples of labeled tweets are presented in Figure 1.
-Figure 2 displays the class distribution in WildFireCan-MMD.
-Also, for training and testing the custom classifiers, given in Section 4, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets.
-The size of labeled classes are presented in Table 1.
-Figure 2: Class distribution of WildFireCan-MMD.
 4 Classification Methodology
-This section describes a range of algorithms evaluated for the 13-class classification of WildFireCan-MMD, utilizing text and image modalities.
-We also experimented with using Optical Character Recognition (OCR) to further extract more information from the images and create another channel of text to aid in classification.
-We experiment with using VLMs as zero-shot classifiers, and compare these results with fine-tuned transformer-based classifiers as well as to baseline models such as Random Forest, and Gaussian Naive Bayes.
-We present the implementation details of classification algorithms in the following.
-4.1 VLM as Zero-Shot Classifiers
-We crafted classification prompts for both gpt-4o-mini and LLaVa in an ’all-at-once’ method, meaning that with each query to the model, a list of all categories and their descriptions are given as a prompt, along with a sample (text + image) from the data to classify.
+This section describes a range of algorithms evaluated for the 13-class (12 categories of the taxonomy +“Other”) classification of WildFireCan-MMD.
+In general, we employ VLMs as zero-shot classifiers, and compare the results with fine-tuned transformer-based classifiers as well as to classical methods such as decision tree, and Gaussian naive bayes, as minimum baselines.
+Since each post on X can have up to 4 associated images, we duplicated the text and paired with each image as an individual sample.
+For training and testing the custom classifiers, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets.
+We use the same test set for all of the classifiers, including zero-shot VLMs.
+4.1 VLMs as Zero-Shot Classifiers:
+We considered three VLMs, GPT-4 Achiam et al.
+(2023), LLaVA Liu et al.
+(2023), and SmolVLM Marafioti et al.
+(2025) and tested them in a zero-shot setting.
+We created classification prompts for all three VLMs where, with each query to the model, a list of all categories and their descriptions is given as a prompt along with a data instance (text + image) to classify.
 We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter.
 The system and user prompts can be seen in Table C.1.
-4.1.1 gpt-4o-mini:
-We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18.
-A specified content field in each query contains our system prompt, text from the Tweet, and user prompt.
-The image for the tweet is also sent with this payload.
-The ’temperature’ is specified as being $0.1$ for reproducibility.
-’max_tokens’ is set to be $1024$, as the text plus the image plus the model’s response will never be greater than this.
-The responses are saved as a JSONL file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1.
-This then allowed us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
-4.1.2 LLaVA:
-As an open source model, querying LLaVA required a different approach than gpt-4o-mini.
-For our work, we opted for the 13B 4-bit version of LLaVa.
-The temperature was again set to $0.1$, and max_tokens set to $1024$.
-A custom conversation template was created using the original LLaVA codebase to construct our system prompt using the roles USER and ASSISTANT.
-The question file was created using the steps discussed above for gpt-4o-mini, the only difference being the addition of ’<>’ characters around each option as recommended by the source [^4] https://github.com/haotian-liu/LLaVA/blob/main/docs/Evaluation.md for querying LLaVA with a multiple-choice prompt.
-The text from the tweet is passed similarly to the method described above for gpt-4o-mini, and the image is converted to a tensor and passed with this to the model.
-The answers are saved similarly as described for gpt-4o-mini and saved, allowing us to calculate metrics such as balanced accuracy score and weighted precision, recall, and F1-score.
+gpt-4o-mini: We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18.
+A specified content field in each query contains our system prompt, text and image from the tweet, and user prompt.
+The responses are saved as a file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1.
+This then allows for the calculation of metrics.
+LLaVA: We opted for the 13B 4-bit version of LLaVA due to hardware constraints.
+Using a combination of a publicly available Colab notebook  [^4] https://github.com/camenduru/LLaVA-colab and the original code from the open-source LLaVA repository  [^5] https://github.com/haotian-liu/LLaVA, the model was loaded onto the GPU within Colab.
+The answers are saved similarly as described for gpt-4o-mini.
+smolVLM: Being the smallest of the VLMs experimented with, only requiring <10GB of GPU ram, experiments were able to be performed locally without quantization.
+Due to its inclusion in the Transformers library, loading and querying the model with the previously created prompts was simpler and cheaper than the other methods, while taking a similar amount of time to process.
 4.2 Custom Classifiers
-To create a custom multimodal classifier for our dataset, the following steps were taken to implement concatenative early and late fusion.
-Early and late fusion 2-head structure::This model consists of two architectures working in parallel, an image head and a text head.
-The image head uses the ViT transformer, and the text head uses the RoBERTa transformer.
-Another linear layer is added for prediction.
+Prior work by Koshy and Elango (2023) demonstrated that a multimodal architecture combining RoBERTa with a BiLSTM and ViT, using multiplicative fusion and an additional attention mechanism, achieved the best performance on the CrisisMMD dataset.
+We evaluated this architecture, among several others, on our WildFireCA-MMD dataset.
+Our findings indicate that the BiLSTM and additional attention components did not improve performance in our context.
+Instead, a simpler architecture employing early fusion via feature concatenation of RoBERTa and ViT embeddings yielded the best results.
+We experimented with several model structures described in the following.
+2-head structure:This model consists of two parallel architectures, an image head and a text head.
+The image head uses the ViT666’google/vit-base-patch16-384’ transformer, and the text head uses the RoBERTa777’roberta-base’ transformer.
+A linear layer is added for prediction.
 Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion.
-We created another model with the same strategy but utilizing late fusion instead, where each head has its own final linear layer and functions as an independent classifier.
-The two heads output probabilities, which are then averaged for a final result.
-Training and Evaluation: A random seed was used for training reproducibility.
-Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by google/vit-base-patch16-384), and a max length for the text of 200.
-Twitter posts can have a maximum length of 200 characters, but after viewing the average and max length of all posts within the dataset, we learned that most users don’t go anywhere near this limit, and so this value could be decreased.
+3-head structure: To enhance classification without more manual labeling, we extracted text from images using EasyOCR [^8] https://github.com/JaidedAI/EasyOCR.
+The extracted text was saved in a new column (extracted_text), defaulting to a filler if none was found.
+Our classifier was updated to include a third RoBERTa head for this column, as shown in Figure D.2.
+Training and Evaluation: A random seed was used for reproducibility.
+Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by our ViT model).
 We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss.
 Our experimentation found that the model converged at around 10-15 epochs.
-The model was then tested on the test set (9̃00 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
-Additional Information with OCR: Trying to improve our classification metrics without gathering and labeling more data, we considered other ways in which we could extract more information from the data we had.
-Something we noticed in much of the data, specifically the Evacuees, Smoke & Air Quality, General Information, Preparedness, Warnings and Status Updates, and Weather Reports classes, was that many of the images contained valuable text which we believed could aid in classification.
-Although ViT and RoBERTA both use the transformer architecture, the way text is processed differs from images and doesn’t allow for ViT to differentiate text within images from the surrounding pixels in this current implementation.
-This led us to consider ways in which we could extract text from the images in our dataset using the EasyOCR tool [^5] https://github.com/JaidedAI/EasyOCR.
-EasyOCR works by taking an image, preprocessing it, then applying the CRAFT algorithm to find bounding-box regions of text in an image.
-The pixels in these regions are sent to a text recognition model composed of a CRNN with three components: feature extraction, sequence labeling, and decoding to extract the text from these regions and turn them into strings as output.
-Deploying this model on our dataset, we extracted all of the text from every image and saved it to a new column as a pre-processing step.
-As a result, our dataset contains four columns: text, extracted_text, images, and labels.
-If an image was found to not have any text, a filler sentence was used: ’Image does not contain text.’.
-Utilizing this extra column of text data, we added another RoBERTa head to our classifier.
-This new classifier, now with three separate heads concatenated and fed into a single linear layer, as appeared in Figure D.2.
-Note that this figure represents the early fusion version.
-A late fusion version was also created similarly as described for the 2-head classifier.
-The model was then trained with the same settings as described previously, for a similar number of epochs.
+The model was then tested on the test set (~1000 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
+4.3 Baseline Classifiers
+To establish baseline, we evaluated several classic classifiers on our dataset, namely Decision Tree, Gaussian Naive Bayes, and K-Nearest Neighbors (KNN).
+For textual data, we extracted TF-IDF feature matrices by applying the TfidfVectorizer with max_features=5000.
+The image data, was resized to 64$\times$64 pixels.
+The resulting TF-IDF and image matrices were concatenated and used as input to the classifiers.
+Hyperparameters were set to default values in Scikit-learn  [^9] https://scikit-learn.org/stable/.
 5 Experimental Results
-Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and two baseline models per class and as an average over all classes.
-Also, figure E.1 shows the confusion matrices of the VLMs and the custom classifiers.
+Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and baseline models per class and as an average over all classes.
+Also, Figures E.1 and E.2 show the confusion matrices of the custom classifiers and VLMs, respectively.
 Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
-Table 2: F1-scores per-class for VLMs, custom classifiers, and baseline models.
+Some classes are harder to distinguish than others.
+In general, we observe a large variance of classification F-scores across classes.
+For example, all classifiers struggle with classifying examples of Evacuees and General Information.
+We explain this observation by the great variability of information in these classes.
+On the other hand, classes with less information variety, such as Insurance and Smoke and Air Quality, are more manageable for most classifiers to detect.
+Table 2: F1-scores per-class for VLMs, custom and baseline classifiers. Maximum results are in bold.
 VLMs Custom Classifiers Baseline Classifiers
-Class LLaVa gpt-4o-mini 2-head-early 3-head-early 2-head-late 3-head-late R. Forest Gaussian NB
-Evacuees 0.17 0.37 0.77 0.73 0.79 0.74 0.29 0.50
-General Information 0.14 0.16 0.59 0.46 0.53 0.59 0.35 0.67
-Preparedness 0.61 0.65 0.82 0.81 0.86 0.84 0.52 0.80
-Weather Reports 0.59 0.64 0.88 0.91 0.91 0.91 0.45 0.66
-Warnings & Status Updates 0.36 0.45 0.80 0.76 0.77 0.74 0.50 0.57
-Reports of Actions of Responders 0.29 0.63 0.89 0.86 0.85 0.83 0.48 0.74
-Infrastructure 0.38 0.40 0.92 0.95 0.90 0.93 0.62 0.64
-Political 0.70 0.77 0.84 0.84 0.84 0.85 0.28 0.78
-Insurance 0.93 0.90 0.93 0.89 0.93 0.91 0.92 0.91
-Advertisement 0.08 0.68 0.76 0.76 0.74 0.75 0.23 0.81
-Smoke and Air Quality 0.76 0.79 0.86 0.86 0.86 0.85 0.58 0.70
-Support 0.25 0.55 0.73 0.63 0.75 0.78 0.45 0.63
-Other 0.05 0.54 0.78 0.78 0.76 0.79 0.36 0.60
-Average 0.41 0.58 0.81 0.79 0.81 0.81 0.46 0.69
-Weighted Average 0.45 0.60 0.83 0.81 0.82 0.82 0.48 0.67
-LLaVa has limited capabilities.Testing on WildFireCan-MMD, We observed that LLaVa is not able to reach an acceptable result, performing close to the Random Forest baseline and underperforming gpt-4o-mini by a large margin.
-Specifically, we noted a large difference in F1-score between gpt-4o-mini and LLaVA for the Advertisement and Other classes (gpt4o-mini achieved between 0.5 and 0.6, whereas LLaVA reaches only less than 0.1).
-This can be explained by the fact that these two classes have the highest variance by far and are hard to handle for LLaVa.
-Specifically, the Other category is very diverse since tweets are labeled as so if they don’t fit into any of the other categories.
-Referring to confusion matrices (Figure E.1(a)), we can see that LLaVA struggles with labeling the Warnings & Status Updates class, where it often classified tweets from other classes as such, especially the Other class.
-gpt-4o-mini shows strengths and limitations across various categories.
-On the other hand, gpt-4o-mini, being much larger and more sophisticated than LLaVA, outperforms it in almost every class except for the Insurance category, where it has been beaten by LLaVA by a couple of percent.
-Interestingly, gpt-4o-mini reaches a similar F1-score as custom classifiers on the Smoke & Air Quality, Advertisement, Insurance and Political classes, which combined encompass over a third of the dataset.
-This is probably due to the lower variance of these classes.
-On the contrary, the classes in which it suffers the most are Evacuees, General Information, Warnings & Status Updates, Infrastructure which contain some of the highest variance.
-It is difficult to construct a prompt, which encompasses the variance of these classes, while staying within the token limit.
-So, this may be one reason why gpt4-o performed worse than custom models in these categories.
-As seen in Figure E.1(b) we can see that gpt-4o-mini had the most trouble with the Warnings & Status Updates class, where it often confused it with the General Information and Evacuees classes.
-This effect caused poor performance in the Infrastructure class as well.
+Class LLaVA GPT-4o-mini smolVLM 2-head 3-head Dec. Tree Gauss. NB KNN
+Evacuees 0.17 0.37 0.25 0.77 0.73 0.47 0.53 0.29
+General Information 0.14 0.16 0.00 0.59 0.46 0.23 0.46 0.21
+Preparedness 0.61 0.65 0.23 0.82 0.81 0.57 0.72 0.49
+Weather Reports 0.59 0.64 0.00 0.88 0.91 0.46 0.63 0.52
+Warnings & Status Updates 0.36 0.45 0.00 0.80 0.76 0.45 0.59 0.43
+Reports of Actions of Responders 0.29 0.63 0.00 0.89 0.86 0.45 0.70 0.40
+Infrastructure 0.38 0.40 0.00 0.92 0.95 0.47 0.75 0.52
+Political 0.70 0.77 0.00 0.84 0.84 0.52 0.75 0.31
+Insurance 0.93 0.90 0.92 0.93 0.89 0.93 0.93 0.89
+Advertisement 0.08 0.68 0.00 0.76 0.76 0.22 0.65 0.40
+Smoke and Air Quality 0.76 0.79 0.50 0.86 0.86 0.60 0.69 0.51
+Support 0.25 0.55 0.06 0.73 0.63 0.30 0.60 0.33
+Other 0.05 0.54 0.27 0.78 0.78 0.35 0.67 0.33
+Weighted Average 0.45 0.60 0.20 0.83 0.81 0.48 0.67 0.43
+VLMs struggle with nuanced disaster communication, underperforming compared to classic classifiers.
+Table 2 shows that VLMs, including LLaVA, GPT-4o-mini, and smolVLM, consistently underperform across most disaster-related classes.
+Their weighted F1-scores range from 0.20 (smolVLM) to 0.60 (GPT-4o-mini), all falling short of the best-performing baseline (Gaussian Naive Bayes, 0.67).
+Similarly, the Decision Tree classifier achieved a weighted score of 0.48, comparable to LLaVA’s 0.45.
+The performance of smolVLM in particular was notably poor, with a weighted average of only 0.20 and zero scores in multiple categories.
+These results suggest that, despite their general-purpose capabilities, current VLMs struggle to perform reliably in domain-specific, fine-grained classification tasks, such as the case for WildFireCan-MMD.
 Custom classifiers are more robust and reliable.
-Referring to Table 2, we can see that the two and three-head classifiers provide similar results.
-Our 2-head classifiers have an average weighted F1-score of 0.83 for early fusion, and 0.82 for late fusion.
-The 3-head classifiers have an average weighted F1-score of 0.81 for early fusion, and 0.82 for late fusion.
-Surprisingly, the three-head classifier performs slightly worse in most classes than the two-head classifier, even though it receives extra contextual information from the images.
-One important consideration, however, when thinking about the utility and purpose of this work, is that the 3-head classifiers perform slightly better on key classes such as Infrastructure, Weather Reports, Smoke & Air Quality, General Information, and Support.
-Infrastructure is an important class because it can provide information to first responders and administrative personnel about key infrastructure that has been damaged or is currently in need of focus.
-Weather Reports is also important because it can provide insights and predictions about weather patterns and likely areas of new fires.
-General Information is useful to administrators and the general public as this class contains broad information about the situation as a whole.
+In contrast to VLMs, our custom classifiers (2-head and 3-head architectures) demonstrate superior and more stable performance.
+With weighted F1-scores of 0.83 and 0.81, they outperform all baselines and VLMs across nearly all categories.
+The 2-head classifier achieves top scores in Insurance (0.93), Reports of Actions of Responders (0.89), Smoke and Air Quality (0.86).
+The 3-head model benefits from incorporating OCR-extracted text, performing particularly well in Infrastructure (0.95), and Weather reports (0.91).
+While incorporating OCR-extracted textual information doesn’t improve the classification results overall, it might be a favorable approach when certain classes are deemed to be more critical for certain response strategies.
+6 Analysis of Trends in Data
+After labeling our dataset and determining our best-performing classifier (the 2-head classifier described in section 4.3 and Figure D.1), we proceeded to collect another 46,279 posts from X.
+The collection methodology was the same as explained in section 3.1, where we gathered tweets using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions.
+We combined this with a keyword-based search, then combined the results and removed duplicates.
+The full set of queries used for collecting this new data can be seen in Appendix F.
+First, we extracted the author locations from the samples.
+Our goal was to obtain the province-level region where each sample is coming from.
+After extracting the author locations, which could be either automatically generated or user-provided, we ran the queries through OSMNX [^10] https://osmnx.readthedocs.io/en/stable/.
+This allowed us to automatically search for the locations of cities or towns that users provided and extract the province level information.
+If a users location was not found, or not in Canada, it was labeled as such.
+The distribution of posts by province can be seen in Figure 2, showing that the majority of data is posted from British Columbia and Alberta.
+Additional information about this data, including a word cloud of the text and the frequency distribution of the posts by word count can be found in Figures 1(a) and 1(b).
+Figure 2: Distribution of collected tweets based on author location.
+Using our best-performing classifier with an F1-score of 83%, we obtained predictions for the new data.
+This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample.
+This of course, will not uphold the same gold standard of labeling that a human annotator could achieve, and manual labeling will have to be done in the future.
+Individual classifications should be approached with skepticism, but this method should still allow for general trends in the data to be analyzed on a broader scale.
+Figure 3 (left) shows the total number of posts per week sorted by year.
+This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country Reuters Staff (2023); CBC News (2024).
+Figure 3 (right) illustrates weekly trends in predicted labels during 2023.
+Sharp spikes in Smoke and Air Quality and Warnings & Status Updates likely correspond to major wildfire events occurring in Alberta and BC.
+Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated.
+Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 Williams (2023).
+Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
+Figure 3: Weekly post distributions: (left) overall yearly trend, (right) class-wise distribution in 2023.
 7 Discussion
@@
 On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management.
-If performance is improved, the open-source nature of this model might make it suitable for disaster response, allowing compliance with responsible AI guidelines through offering access to the code base and ensuring ethical scrutiny.
-Particularly, the transparent nature of open source is crucial when handling sensitive data that may include personal information.
-Despite gpt-4o-mini performing better on our test set, it is important to keep these above considerations in mind.
-When labeled data is unavailable, VLMs can function as a zero-shot classifier, but in practice, neither model is particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
+Also, smolVLM, while being both open-source and computationally affordable, has a performance far below acceptable rates.
+Therefore, when labeled data is unavailable, VLMs might be considered as a zero-shot classifier, but in practice, these models are not particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
 Fine-tuning on task-specific data is essential.
-Results presented in Table 2 demonstrate that even simple approaches such as Random Forest and Gaussian NB, when trained with task-specific labelled data, provide similar or better results than what is achieved from zero-shot VLMs.
+Results presented in Table 2 demonstrate that even simple approaches such as decision tree and Gaussian NB, when trained with task-specific labeled data, provide similar or better results than what is achieved from zero-shot VLMs.
 Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier.
-Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning advanced classification structures, such as pre-trained transformer-based text and image encoders.
-The custom classifiers we have developed serve as a foundational step for triaging tweets during wildfires, providing an initial performance benchmark and demonstrating our approach’s feasibility.
-However, to enhance the accuracy and robustness of our model, future work will need to adopt a more comprehensive approach to experimentation.
-One area for improvement is the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
+Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning common classification structures, such as pre-trained transformer-based text and image encoders.
+These findings align with other works such as Nadeem et al.
+(2025).
+Custom classifiers can be used to discover social media use patterns.
+The primary function of the custom classifier is to perform real-time labeling of social media data during wildfire events, enabling the timely extraction of actionable information to support disaster response efforts.
+Beyond real-time deployment, our post-hoc analysis of a large unlabeled dataset illustrates the classifier’s capacity to identify latent patterns and temporal trends, providing a scalable tool for retrospective analysis and evidence-based policy development.
+To enhance the accuracy and robustness of our model, future work can consider the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
+Fine-tuning of VLMs is also a promising avenue of improving the task which should be explored in future works Duan et al.
+(2024); Zhai et al.
+(2024).
 8 Conclusion
-General-purpose VLMs, while powerful, struggle to perform reliably in high-stakes, context-rich scenarios like disaster response.
-To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts (text + image) curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation.
-Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs.
+Despite the advancements in general-purpose VLMs, they struggle to perform reliably in high-stakes, context-rich scenarios like disaster response.
+This low accuracy, alongside the resource-intensive nature of such models, necessitates developing smaller fine-tuned models, specialized for specific regions and contexts.
+To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation.
+Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs, and suitable for training specialized models.
 As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
+9 Acknowledgment
+This work was funded by the National Program Office of the National Research Council Canada, under the New Beginning Initiative Program.
 Limitations
 While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
-The fisrt limitation is the scope and size of the dataset.
-Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate general-purpose models.
+The first limitation is the scope and size of the dataset.
+Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models.
 This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
@@
 Moreover, zero-shot evaluation of VLMs introduces limitations. Prompt engineering for zero-shot setups is a heuristic process, and better prompts might yield different results.
+Finally, there is a broader risk of misapplication.
+Deploying these models in real-world disaster contexts without appropriate oversight, interpretability mechanisms, or human-in-the-loop verification may lead to misclassification, misinformation amplification, or missed signals, potentially affecting emergency coordination and public trust.
 Appendix A Data collection Queries
-Table A.1 shows the full queries used to collect data from X.
+Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
 Year Query
@@
 2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
-Table A.1: The full queries used for scraping data from X.
+Table A.1: The queries used for collecting the data for WildFireCan-MMD.
 Appendix B Topic modeling parameters
-In order to optimize the performance of BERTopic on our dataset, careful consideration of hyperparameters is key McInnes et al.
+In order to optimize the performance of BERTopic, careful consideration of hyperparameters is key McInnes et al.
 (2018).
@@
 For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
-Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting a set number of topics to sort the data into, as at this stage, we were unsure of what sort of topics were contained in our data.
-BERTopic had sorted our data into over 100 fine-grained topics.
-This was still relatively large to manage, and it was difficult to determine a humanitarian category for each fine-grained topic.
-After some experimentation, we systematically decreased the total number of topics to determine which topics remained over time.
 Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics.
@@
 Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
-Using BERTopic in the way described proved to be useful, as it gave us a good understanding from the beginning of what our future categories were to look like.
-Relying on this method alone was not enough, however, and having to look through the data manually at some point was inevitable.
-Many rounds of manual annotation ensued on a small portion of the data, after having it randomly mixed, in order to avoid constructing our categories on themes that may only exist at a certain time during the wildfire event.
-Through performing these steps, keeping in mind our discovered topics through BERTopic and humanitarian goals, new categories were added or dropped based on what was seen in the data.
-A final round of labelling then commenced, and the entire dataset was manually labeled.
 Keywords: [crews, safe, poles, damaged, continue, lines, working restore, power, conditions, restore]
@@
 Appendix C Prompts for zero-shot classification
-We show the full prompt used to evaluate VLMs in Table C.1.
-Table C.1: System and User prompt used for gpt-4o-mini and LLaVa.
+We show the full prompt used to evaluate the aforementioned VLMs in Table C.1. The following hyperparameters are used for all VLMs: temperature=0.1, max_new_tokens=1024.
+Table C.1: System and User prompt used for VLMs.
 System Prompt:
-"You are an assistant who is being given an image and text pair as a Twitter post which was created during a natural disaster event.
+"You are an assistant who is being given an image and text pair as a Twitter post, which was created during a natural disaster event.
 Your task is to use information from both the text and the image to decide which option the post should be labeled as.
@@
 Figure D.2: Custom multimodal 3-head early fusion classifier structure.
-Appendix E Confusion matrix for all classifiers
-For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figure E.1.
-a) LLaVA, b) gpt-4o-mini, c) 2head-early, d) 3head-early, e) 2head-late, f) 3head-late
-Figure E.1: Confusion matrices of classifiers and VLMs.
+Appendix E Confusion matrices of classifiers
+For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figures E.1 and E.2.
+(a) 2-head classifier.
+(b) 3-head classifier.
+Figure E.1: Confusion matrices of custom classifiers.
+(a) LLaVA.
+(b) GPT-4.
+(c) SmolVLM.
+Figure E.2: Confusion matrices of VLMs.
+Appendix F Queries Used for Additional Data Collection :
+Below is the list of queries and keywords used to collect additional unlabeled data discussed in Section 6.
+2018
+Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire)
+Keyword Query: (alberta OR british columbia OR Prince George OR Grande Praire OR Waterton OR Bulkley Nechako OR Nadina Lake OR Kootenay OR Crowsnest Pass OR Medicine Lake OR Comstock Lake OR Tugwell Creek OR Sooke OR Nanaimo Lakes OR Tweedsmuir OR Johnny Creek OR Alkali Lake OR Lutz Creek OR Shovel Lake OR Nadina Lake OR Verdun Mountain OR Silver Lake OR Tommy Lakes OR Island Lake OR Chutanli Lake) (wildfire OR forest fire)
+2019
+Hashtag Query: (#ABWildfire OR #albertawildfire OR #ABFire)
+Keyword Query: (alberta OR calgary OR edson OR Fort McMurray OR Grande Prairie OR High Level OR Lac La Biche OR Whitecourt OR Steen River OR Chuckegg Creek OR Peace River OR Slave Lake OR Wood Buffalo National Park) (wildfire OR forest fire)
+2020
+Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire)
+Keyword Query: (british columbia OR alberta OR sasksatchewan OR yukon OR northwest territories OR manitoba OR ontario OR quebec) (wildfire OR forest fire)
+2021
+Hashtag Query: (#MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #pafire OR #ontariofire OR #manitobafire OR #sasksatchewanfire)
+Keyword Query: (manitoba OR ontario OR sasksatchewan OR british columbia) (wildfire OR forest fire)
+2022
+Hashtag Query: (#YTwildfire OR #yukonwildfire OR #YTfire OR #yukonforestfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #nwtforestfire)
+Keyword Query: (#yzf OR #nwt OR #Yellowknife OR Yukon OR Northwest Territories OR Whitehorse OR Yellowknife OR Dawson City OR Great Slave Lake OR Norman Wells OR Inuvik OR Watson Lake OR Hay River OR Fort Smith OR Tuktoyaktuk OR Behchoko) (wildfire OR forest fire) (-alaska -Eielson -CityofNorthPole)
+2023
+Hashtag Query: (#BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #SKwildfire OR #sasksatchewanwildfire OR #SKfire OR #YTwildfire OR #yukonwildfire OR #YTfire OR #NTwildfire OR #northwestterritorieswildfire OR #NTfire OR #NWTwildfire OR #NWTfire OR #MBwildfire OR #manitobawildfire OR #MBfire OR #ONwildfire OR #ontariowildfire OR #QCwildfire OR #quebecwildfire OR #QCfire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
+Keyword Query: (ontario OR quebec OR sasksatchewan OR british columbia OR manitoba OR northwest territories OR yukon OR alberta) (wildfire OR forest fire)
+2024
+Hashtag Query: (#JasperStrong OR #JasperWildfire OR #JasperAB OR #BCwildfire OR #britishcolumbiawildfire OR #BCfire OR #ABwildfire OR #albertawildfire OR #ABFire OR #CanadaOnFire OR #CanadaWildfire OR #CanadaFires OR #CanadaIsOnFire)
+Keyword Query: (canada OR ontario OR quebec OR sasksatchewan OR manitoba OR northwest territories OR yukon OR british columbia OR alberta OR jasper) (wildfire OR forest fire)
+Note: Every query above includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
+Appendix G Unlabeled dataset
+A word cloud of the text data of the unlabeled dataset collected in Section 6, and also the frequency distribution of the posts by word count, are shown in Figures 1(a) and 1(b).
+(a) Word cloud of the posts.
+(b) Frquency distribution of the posts w.r.t their lengths.
+Figure G.1: Word cloud of the textual part of the unlabeled tweets and the frequency distribution of them regarding their word counts.

### v3

@@
 Social media offers real-time updates, but extracting relevant insights remains a challenge.
+In this work, we focus on multimodal wildfire social media data, which, although existing in current datasets, is currently underrepresented in Canadian contexts.
 We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes.
-Evaluating both vision-language models and custom-trained classifiers, we show that while zero-shot prompting offers quick deployment, even simple trained models outperform them when labelled data is available.
-Our best-performing transformer-based fine-tuned model reaches 83% f-score, outperforming gpt4 by 23%.
-Also, we demonstrate how this model can be used to uncover trends during wildfires.
-Our findings highlight the enduring importance of tailored datasets and task-specific training.
+We evaluate zero-shot vision-language models on this dataset and compare their results with those of custom-trained and baseline classifiers.
+We show that while baseline methods and zero-shot prompting offer quick deployment, custom-trained models outperform them when labelled data is available.
+Our best-performing custom model reaches 84.48±0.69% f-score, outperforming VLMs and baseline classifiers.
+We also demonstrate how this model can be used to uncover trends during wildfires, through the collection and analysis of a large unlabeled dataset.
+Our dataset facilitates future research in wildfire response, and our findings highlight the importance of tailored datasets and task-specific training.
 Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
 WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada
-Braeden Sherritt Carleton University                        Isar Nejadgholi National Research Council Canada                        Marzieh Amini Carleton University
+Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,MarziehAmini@cunet.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca
 1 Introduction
@@
 As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked.
-Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising about %12 of the total dataset and specific to the US context.
-The HumAID dataset by Alam et al.
-(2021) also contains wildfire data from Canada, but only a couple of thousand samples from 2016.
+Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising 1589 samples, and is specific to the California wildfires of October 2017.
+The wildfire-related samples from Canada in the HumAID dataset Alam et al.
+(2021) are limited to a single month in 2016, as well as another month of California samples from 2018.
 Therefore, wildfires in Canada remain underexplored in social media analysis research.
@@
 Using our developed taxonomy, we manually labeled 4,688 wildfire-related tweets from Canada and created WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires.
-To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several VLMs, as well as the feasibility of training custom classifiers.
-Our best-performing classifier is a fine-tuned model, which uses RoBERTa Liu (2019) and Vision Transformer (ViT) models Ofli et al.
-(2020); Basit et al.
-(2023) to jointly process text and image parts of tweets.
-While WildFireCan-MMD is sourced from data posted in 2022-2024, we collected an additional set of unlabeled tweets across seven years (2018-2025).
-We then used our best-performing classifier to analyze this data and uncover trends and patterns of social media use in Canada.
-The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at WildFireCanMMD-Github [^1] Please contact the authors for more information about data sharability., under CC-BY-NC license.
-The contributions of this study are summarized as follows:
-  • We devised a twelve-category taxonomy through multimodal topic modeling of Canadian wildfire-related user-generated content, and introduced WildFireCan-MMD, the first Canada-specific multimodal wildfire dataset, comprising 4,688 posts from X in 2022–2024, manually labeled based on our taxonomy.
-  • We compared custom multimodal transformer models trained on WildFireCan-MMD with several VLMs and classical methods.
-Fine-tuned classifiers outperform zero-shot VLMs significantly; our best model achieves an F-score of 0.83, exceeding gpt-4o-mini by 23%.
-  • We collected an additional 46,279 posts from X, spanning across seven years, used our trained classifier to label them, and reported the uncovered trends and patterns of social media use during wildfires in Canada.
+To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several Vision-Language Models (VLMs), custom deep-learning classifiers, as well as classical approaches, as a minimum achievable performance.
+Our best-performing classifier utilizes RoBERTa Liu (2019) and Vision Transformer (ViT) models to jointly process the text and image components of tweets.
+A custom transformer-based fusion module is designed to enable the modalities to interact deeply.
+While WildFireCan-MMD is sourced from data posted between 2022 and 2024, we collected an additional set of unlabeled tweets spanning seven years (2018-2024).
+Using our best-performing classifier, we analyze this unlabeled data and uncover trends and patterns of social media use during wildfires in Canada.
+The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at [Anonymized GitHub], under CC-BY License.
+The contributions of this study are summarized as:
+  • We collected 4,688 wildfire-related posts from X in Canada from 2022 to 2024 and devised a twelve-category taxonomy using unsupervised topic modeling and human analysis.
+We introduce WildFireCan-MMD, the first Canada-specific multimodal wildfire social-media dataset, annotated by three annotators.
+  • We propose a custom multimodal transformer model, trained on WildFireCan-MMD and compare it with several VLMs and classical methods.
+The fine-tuned custom classifier outperforms zero-shot VLMs significantly; our model achieves an f-Score of 84.48±0.69%.
+  • Additionally, we collected 46,279 posts from X from 2018 to 2024.
+We used our trained classifier to label them and reported the trends and patterns of social media use during wildfires in Canada.
+Results show that retrospective analysis using a classifier can uncover trends which relate to real situational updates.
 2 Background and Literature Review
@@
 (2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
-Our work is motivated by this literature and focuses on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
+Our work is motivated by this literature as well as the literature shown in Appendix G which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
 2.2 Multimodal classification for disaster response
@@
 (2025).
-Alam et al.
-(2018) addressed the shortage of multimodal labeled social media data for disaster response by introducing CrisisMMD, a widely used dataset containing annotated tweets and images from various natural disasters.
-Later extended by datasets like HumAID Alam et al.
-(2021), CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images.
-It remains the most prominent dataset in this domain, used in many recent studies Basit et al.
-(2023); Koshy and Elango (2023); Shetty et al.
+The study by Alam et al.
+(2018) released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes).
+CrisisMMD contains thousands of tweets labeled for categories like “infrastructure and utility damage” or “rescue volunteering or donation effort”.
+Additionally, Alam et al.
+(2018) later released the HumAID Alam et al.
+(2021) dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019).
+It is similarly annotated for emergency response categories (e.g.
+casualties, infrastructure damage).
+There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex Olteanu et al.
+(2015) and CrisisMMD Alam et al.
+(2018) datasets Wu et al.
 (2022).
-While previous works such as CrisisMMD used a generalized scheme of annotations across various types of different disasters, our work develops a wildfire-specific taxonomy of twelve categories and introduces WildFireCan-MMD, a new Canadian wildfire dataset.
+CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images.
+It remains the most prominent dataset in this domain, used in many recent studies Ofli et al.
+(2020); Khattar and Quadri (2022); Basit et al.
+(2023); Koshy and Elango (2023); Rezk et al.
+(2023) and many others.
+Therefore, there is a notable gap in the in the availability of multimodal wildfire datasets.
+In the context of Canada, where the most common and destructive natural disaster is wildfires, CrisisMMD is lacking in data.
+Although it contains  18k samples from other disasters such as hurricanes, earthquakes and floods, the amount of wildfire data is proportionally small (1589 samples).
+Additionally, this data is only from the California wildfires of October 2017.
+The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016, as well as another month of California samples from 2018.
+Therefore, wildfires in Canada remain underexplored in social media analysis research.
+To address this limitation, this work focuses on gathering social media data related to Canadian wildfires, and proposes WildFireCan-MMD, a new Canadian wildfire social-media dataset with a wildfire-specific taxonomy 12 categories.
 Algorithms: Previous work by Ofli et al.
-(2020) introduced a joint representation learning from two parallel deep learning architectures where one architecture represents the text modality and the other represents the image modality.
-Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures, respectively, followed by a dense layer for prediction.
+(2020) introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality.
+Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures respectively, followed by a dense layer for prediction.
 The work by Basit et al.
@@
 (2018).
-We build on the previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
+Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix G.
+Most previous studies use a CNN for the image modality and then a BERT-based transformer for the text modality.
+Other works sometimes use ViT for the image modality.
+Concatenative fusion of the two modalities into a series of linear layers seems to be the most common approach, with other works using multiplicative or additive fusion.
+Cross-attention fusion of the two modalities seems to be another popular approach that seems to give promising results.
+The approach given by Pranesh (2022) proposes a transformer-like structure to fuse the two modalities.
+We build on the this previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
 In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al.
@@
 3.1 Data Collection
-To construct our dataset, we utilized X’s Pro-Tier paid API to collect a diverse set of posts surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024.
-The dataset comprises 4,688 unique image-text pairs gathered from tweets posted between April and June of each respective year.
-Data collection was conducted using a hashtag-based search strategy.
-Table A.1 presents the full queries used for data collection.
-To curate the hashtags, we started with seed hashtags and used an iterative process of adding new hashtags by collecting sample tweets.
-This approach was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location.
-Instead, we identified and utilized hashtags that, through iterative experimentation, were found to be effective indicators of relevant content originating from Canada during wildfire events.
+To construct the dataset, X’s Pro-Tier paid API was utilized to collect a diverse set of posts, including images and text, surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024.
+The dataset comprises 4,688 unique image-text pairs gathered from the wildfire season (April-September) of each respective year.
+Data collection was conducted using a hashtag-based search strategy as seen in other works Alam et al.
+(2018); Gupta and Hewett (2020); Cruickshank and Carley (2020).
+Appendix A presents the full queries used for data collection.
+To curate the hashtags, basic hashtags such as "#wildfire, #forestfire" were first used.
+Then, through an iterative process of adding new hashtags by collecting sample tweets, more relevant hashtags were discovered.
+This approach of using hashtags was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location.
+Instead, hashtags identified through iterative experimentation were found to be effective indicators of relevant content originating from Canada during wildfire events.
 The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions.
-The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region or county from which the tweet originated.
+The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region from which the tweet originated.
 3.2 Taxonomy Development
-To examine the content of the tweets and discover the hidden latent semantics of our data, we used BERTopic to perform multimodal unsupervised topic modeling Grootendorst (2022); Falkenberg et al.
-(2022); Egger and Yu (2022); Medvecki et al.
-(2023).
-Further explanation of the BERTopic clustering method are provided in Appendix B.
-Employing BERTopic on our dataset, we first let the algorithm determine the number of latent topics instead of setting number of topics to sort the data into.
-When the number of clusters were unspecified, BERTopic had created over 100 fine-grained topics.
-This was relatively large to manage, and difficult to determine a humanitarian category for each fine-grained topic.
-We gradually decreased the number of topics, and for each number, we manually inspected the representative visual samples of topics and keywords generated by BERTopic.
-Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics.
-Some examples of useful topics were posts about infrastructure, weather reports, and reports of actions of first responders.
-We also noticed some less useful posts, such as advertisements and political debates/complaints.
-Some examples of these topics, including representative images and corresponding keywords, generated by BERTopic, can be seen in Figure B.1.
-We then manually inspected the topics and selected 12 classes that were informative and consistently appeared across a varying maximum number of clusters to be considered designated categories for this dataset.
-The full list of categories and their descriptions is given in Table 1.
-We also include an “Other” to accommodate the examples that do not fit into any of these 12 categories.
-The selection of classes was guided by two primary considerations: the frequency of occurrence observed through manual annotation and the perceived relevance of classes to stakeholders in disaster scenarios.
+To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst (2022).
+Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic.
+The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15.
+Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords.
+Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics.
+Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased.
+For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected.
+Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear.
+Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders.
+Less useful posts were also noticed, such as advertisements and political debates/complaints.
+Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix C.
+The topics were inspected and 12 classes were decided upon.
+The full list of classes and their descriptions is given in Table 1.
+To ensure the reliability of the labeling process, the definitions and boundaries of each category were discussed among the research team and a shared understanding was reached.
+Through this discussion, it was determined that the classes are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image).
+This discussion, along with unsupervised topic modeling (BERTopic) were used to find strong topics which were then later assigned to a class.
+Also, the selection of classes was guided by the topics created by BERTopic, observations made through manual annotation, and the perceived relevance of classes to stakeholders in disaster scenarios.
 For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel.
@@
 The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
+An ‘Other’ class is added to accommodate the samples that do not fit into any of the 12 classes.
 Table 1: Categories of WildFireCan-MMD, their descriptions and total count of each class.
 Category Description Count
-Evacuees Information relating to evacuees, their movements, needs, location. 242
-Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1070
-General Information General info about the wildfire situation, such as total hectares burned. 186
-Preparedness Information for the general public to prepare themselves and their property. 265
-Weather Reports Information relating to the weather with a specific location mentioned. 316
-Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 620
-Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 378
-Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 293
-Political Posts directed towards political figures or parties (excluding situation updates). 322
-Support Mental health and financial support, temporary housing for livestock. 164
-Insurance Information relating to insurance, employment insurance, and EI benefits. 142
-Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 116
-Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 569
+Evacuees Information relating to evacuees, their movements, needs, location. 252
+Smoke and Air Quality Tweets related to or showing signs of smoke or the current air quality. 1128
+General Information General info about the wildfire situation, such as total hectares burned. 170
+Preparedness Information for the general public to prepare themselves and their property. 264
+Weather Reports Information relating to the weather with a specific location mentioned. 296
+Warnings & Status Updates Fire bans in certain areas, new information about a specific area, updates from officials. 669
+Reports of Actions of Responders Actions of responders within specific areas or times, including prescribed burns. 356
+Infrastructure Detours, road closures, damage to infrastructure (e.g., utility poles, highways), repairs by crews. 264
+Political Posts directed towards political figures or parties (excluding situation updates). 329
+Support Mental health and financial support, temporary housing for livestock. 178
+Insurance Information relating to insurance, employment insurance, and EI benefits. 158
+Advertisement Posts about food, restaurants, off-topic ads for services (e.g., apps, air purifiers, not insurance-related). 117
+Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 507
 3.3 Data Labeling
-We labeled the collected dataset according to the taxonomy of categories presented in Table 1.
-Over the course of labeling, we ensured that the data did not contain personal information.
-Labeling was carried out in-house by one of the authors using Label Studio [^2] https://labelstud.io/.
-To ensure the reliability of the labeling process, we first discussed the definitions and boundaries of each category among the research team and reached a shared understanding.
-Through this discussion, we determined that the task was not largely subjective, as the categories are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image).
-Therefore, we did not consider multi-annotator labeling necessary.
+Labeling was carried out in-house by three independent annotators using Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 1.
 Examples of labeled tweets are presented in Figure 1.
-The size of labeled classes are presented in Table 1.
+Each image-text pair was given a single label.
+In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample.
+For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’.
+In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone.
+If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided.
+The size of labeled classes is presented in Table 1.
+Annotation was done by three independent annotators, all authors of this paper.
+The annotators are male, 20-30 year old, fluent in English and reside in Canada.
+First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes.
+Over the course of labeling, samples were flagged that contained personal information.
+A vote between the labels from the three annotators decided the final labels.
+Annotation agreement scores can be seen in Table 2.
+A majority agreement was reached on 88.4% of the labels.
+Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels.
+From the results, annotators 1 & 2 had the best agreement scores.
+This is further demonstrated by Cohen’s and Fleiss’ Kappa metrics, which indicate that a moderate-to-good level of agreement is achieved between the independent annotators.
+In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators.
+Additional statistics can be found in Appendix B.
+Table 2: Annotation agreement statistics.
+Metric Agreement
+Majority Agreement (2 same) 88.4%
+Full Agreement (all same) 45.1%
+Vote between all/annotator 1 78.7%
+Vote between all/annotator 2 77.7%
+Vote between all/annotator 3 65.8%
+Cohen’s Kappa annotator 1-2 63.5%
+Cohen’s Kappa annotator 1-3 50.6%
+Cohen’s Kappa annotator 2-3 49.2%
+Fleiss’ Kappa 54.3%
 Label: Reports of Actions of Responders
@@
 4 Classification Methodology
-This section describes a range of algorithms evaluated for the 13-class (12 categories of the taxonomy +“Other”) classification of WildFireCan-MMD.
-In general, we employ VLMs as zero-shot classifiers, and compare the results with fine-tuned transformer-based classifiers as well as to classical methods such as decision tree, and Gaussian naive bayes, as minimum baselines.
-Since each post on X can have up to 4 associated images, we duplicated the text and paired with each image as an individual sample.
-For training and testing the custom classifiers, we split the data into an 80/20 train/test split using stratified sampling, to ensure similar distributions of classes in train and test subsets.
-We use the same test set for all of the classifiers, including zero-shot VLMs.
-4.1 VLMs as Zero-Shot Classifiers:
-We considered three VLMs, GPT-4 Achiam et al.
-(2023), LLaVA Liu et al.
-(2023), and SmolVLM Marafioti et al.
-(2025) and tested them in a zero-shot setting.
-We created classification prompts for all three VLMs where, with each query to the model, a list of all categories and their descriptions is given as a prompt along with a data instance (text + image) to classify.
-We specify that the model’s task is to choose one of the provided categories and respond only with its corresponding letter.
-The system and user prompts can be seen in Table C.1.
-gpt-4o-mini: We used the available API for querying gpt-4o-mini 333version:gpt-4o-mini-2024-07-18.
-A specified content field in each query contains our system prompt, text and image from the tweet, and user prompt.
-The responses are saved as a file where each line contains the question ID from the questions file, as well as the model’s answer supplied as a single letter from the options shown in Table C.1.
-This then allows for the calculation of metrics.
-LLaVA: We opted for the 13B 4-bit version of LLaVA due to hardware constraints.
-Using a combination of a publicly available Colab notebook  [^4] https://github.com/camenduru/LLaVA-colab and the original code from the open-source LLaVA repository  [^5] https://github.com/haotian-liu/LLaVA, the model was loaded onto the GPU within Colab.
-The answers are saved similarly as described for gpt-4o-mini.
-smolVLM: Being the smallest of the VLMs experimented with, only requiring <10GB of GPU ram, experiments were able to be performed locally without quantization.
-Due to its inclusion in the Transformers library, loading and querying the model with the previously created prompts was simpler and cheaper than the other methods, while taking a similar amount of time to process.
-4.2 Custom Classifiers
-Prior work by Koshy and Elango (2023) demonstrated that a multimodal architecture combining RoBERTa with a BiLSTM and ViT, using multiplicative fusion and an additional attention mechanism, achieved the best performance on the CrisisMMD dataset.
-We evaluated this architecture, among several others, on our WildFireCA-MMD dataset.
-Our findings indicate that the BiLSTM and additional attention components did not improve performance in our context.
-Instead, a simpler architecture employing early fusion via feature concatenation of RoBERTa and ViT embeddings yielded the best results.
-We experimented with several model structures described in the following.
-2-head structure:This model consists of two parallel architectures, an image head and a text head.
-The image head uses the ViT666’google/vit-base-patch16-384’ transformer, and the text head uses the RoBERTa777’roberta-base’ transformer.
-A linear layer is added for prediction.
-Figure D.1 shows the structure of our text+image ’2-head’ multimodal classifier using concatenative early fusion.
-3-head structure: To enhance classification without more manual labeling, we extracted text from images using EasyOCR [^8] https://github.com/JaidedAI/EasyOCR.
-The extracted text was saved in a new column (extracted_text), defaulting to a filler if none was found.
-Our classifier was updated to include a third RoBERTa head for this column, as shown in Figure D.2.
-Training and Evaluation: A random seed was used for reproducibility.
-Some notable hyperparameters used include a batch size of 16, an image size of ($384\times{384}\times{3}$) (as required by our ViT model).
-We set the maximum number of epochs to be 20, and used Adam optimizer with a learning rate of $5e^{-5}$, and cross-entropy loss.
-Our experimentation found that the model converged at around 10-15 epochs.
-The model was then tested on the test set (~1000 samples), and metrics were calculated, including balanced accuracy score and weighted precision, recall, and F1-score.
-4.3 Baseline Classifiers
-To establish baseline, we evaluated several classic classifiers on our dataset, namely Decision Tree, Gaussian Naive Bayes, and K-Nearest Neighbors (KNN).
-For textual data, we extracted TF-IDF feature matrices by applying the TfidfVectorizer with max_features=5000.
-The image data, was resized to 64$\times$64 pixels.
-The resulting TF-IDF and image matrices were concatenated and used as input to the classifiers.
-Hyperparameters were set to default values in Scikit-learn  [^9] https://scikit-learn.org/stable/.
+This section describes a range of classifiers evaluated for the 13-class (12 categories + ‘Other’) classification of WildFireCan-MMD.
+We investigate three categories of classification methods.
+First, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers.
+This is the right choice when no training data is available.
+Then we examined classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) to investigate a case where labeled data is available but computation is limited.
+Finally, we experimented with a fine-tuned transformer-based classifier to assess its performance in the presence of labeled data and moderate computational resources.
+Further information on classical methods and VLMs can be found in Appendices E and F.
+Architecture of custom classifier: To choose the architecture of the custom classifier, we experimented with the architectures proposed by various authors on the CrisisMMD dataset (presented in Appendix G) and evaluated them on the WildFireCAD-MMD dataset.
+These included CNNs and transformers for the image modality, as well as various BERT-based transformers for the text modality.
+After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in literature and included in Table G.1 were tested to improve the classification.
+More information about these experiments, as well as an ablation study, can be found in Appendix H.
+The best performing classifier consisted of two parallel encoders, an image and text encoder.
+The image encoder is ViT333’google/vit-base-patch16-384’, and the text encoder is RoBERTa444’roberta-base’.
+The output CLS token of the last hidden state from each are sent through separate projection layers to reduce the dimensionality from 768 to 512.
+The output is passed to the transformer fusion module consisting of two encoder layers, each with d_model=512, nhead=8, dim_feedforward=2048, dropout=0.2.
+A final classification head projects the dimension from 512 to 256, which is followed by ReLU activation and dropout=0.2, then finally from 256 to num_classes.
+Figure I.1 shows the structure of this architecture.
+Training and Evaluation: Three-fold cross validation with stratified sampling was used to train the model and report the results (mean $\pm$ std).
+The random seeds from splitting the data were kept for reproducibility.
+The hyperparameters used were a batch size of 8 and an image size of ($384\times{384}\times{3}$).
+Text sequences are padded to a max length of 144.
+Adam optimizer is used with a learning rate of $1e^{-5}$, weight_decay=0.01, and cross-entropy loss function.
+The model is fully trained after 10 epochs.
+The best model is saved based on f-weighted-average.
+The model was tested on the held-out test set (938 samples).
 5 Experimental Results
-Table 2 shows the calculated F1-scores for the VLMs, custom classifiers, and baseline models per class and as an average over all classes.
-Also, Figures E.1 and E.2 show the confusion matrices of the custom classifiers and VLMs, respectively.
-Note that all metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in our dataset.
-Some classes are harder to distinguish than others.
-In general, we observe a large variance of classification F-scores across classes.
-For example, all classifiers struggle with classifying examples of Evacuees and General Information.
-We explain this observation by the great variability of information in these classes.
-On the other hand, classes with less information variety, such as Insurance and Smoke and Air Quality, are more manageable for most classifiers to detect.
-Table 2: F1-scores per-class for VLMs, custom and baseline classifiers. Maximum results are in bold.
-VLMs Custom Classifiers Baseline Classifiers
-Class LLaVA GPT-4o-mini smolVLM 2-head 3-head Dec. Tree Gauss. NB KNN
-Evacuees 0.17 0.37 0.25 0.77 0.73 0.47 0.53 0.29
-General Information 0.14 0.16 0.00 0.59 0.46 0.23 0.46 0.21
-Preparedness 0.61 0.65 0.23 0.82 0.81 0.57 0.72 0.49
-Weather Reports 0.59 0.64 0.00 0.88 0.91 0.46 0.63 0.52
-Warnings & Status Updates 0.36 0.45 0.00 0.80 0.76 0.45 0.59 0.43
-Reports of Actions of Responders 0.29 0.63 0.00 0.89 0.86 0.45 0.70 0.40
-Infrastructure 0.38 0.40 0.00 0.92 0.95 0.47 0.75 0.52
-Political 0.70 0.77 0.00 0.84 0.84 0.52 0.75 0.31
-Insurance 0.93 0.90 0.92 0.93 0.89 0.93 0.93 0.89
-Advertisement 0.08 0.68 0.00 0.76 0.76 0.22 0.65 0.40
-Smoke and Air Quality 0.76 0.79 0.50 0.86 0.86 0.60 0.69 0.51
-Support 0.25 0.55 0.06 0.73 0.63 0.30 0.60 0.33
-Other 0.05 0.54 0.27 0.78 0.78 0.35 0.67 0.33
-Weighted Average 0.45 0.60 0.20 0.83 0.81 0.48 0.67 0.43
-VLMs struggle with nuanced disaster communication, underperforming compared to classic classifiers.
-Table 2 shows that VLMs, including LLaVA, GPT-4o-mini, and smolVLM, consistently underperform across most disaster-related classes.
-Their weighted F1-scores range from 0.20 (smolVLM) to 0.60 (GPT-4o-mini), all falling short of the best-performing baseline (Gaussian Naive Bayes, 0.67).
-Similarly, the Decision Tree classifier achieved a weighted score of 0.48, comparable to LLaVA’s 0.45.
-The performance of smolVLM in particular was notably poor, with a weighted average of only 0.20 and zero scores in multiple categories.
-These results suggest that, despite their general-purpose capabilities, current VLMs struggle to perform reliably in domain-specific, fine-grained classification tasks, such as the case for WildFireCan-MMD.
-Custom classifiers are more robust and reliable.
-In contrast to VLMs, our custom classifiers (2-head and 3-head architectures) demonstrate superior and more stable performance.
-With weighted F1-scores of 0.83 and 0.81, they outperform all baselines and VLMs across nearly all categories.
-The 2-head classifier achieves top scores in Insurance (0.93), Reports of Actions of Responders (0.89), Smoke and Air Quality (0.86).
-The 3-head model benefits from incorporating OCR-extracted text, performing particularly well in Infrastructure (0.95), and Weather reports (0.91).
-While incorporating OCR-extracted textual information doesn’t improve the classification results overall, it might be a favorable approach when certain classes are deemed to be more critical for certain response strategies.
-6 Analysis of Trends in Data
-After labeling our dataset and determining our best-performing classifier (the 2-head classifier described in section 4.3 and Figure D.1), we proceeded to collect another 46,279 posts from X.
-The collection methodology was the same as explained in section 3.1, where we gathered tweets using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions.
-We combined this with a keyword-based search, then combined the results and removed duplicates.
-The full set of queries used for collecting this new data can be seen in Appendix F.
-First, we extracted the author locations from the samples.
-Our goal was to obtain the province-level region where each sample is coming from.
-After extracting the author locations, which could be either automatically generated or user-provided, we ran the queries through OSMNX [^10] https://osmnx.readthedocs.io/en/stable/.
-This allowed us to automatically search for the locations of cities or towns that users provided and extract the province level information.
-If a users location was not found, or not in Canada, it was labeled as such.
-The distribution of posts by province can be seen in Figure 2, showing that the majority of data is posted from British Columbia and Alberta.
-Additional information about this data, including a word cloud of the text and the frequency distribution of the posts by word count can be found in Figures 1(a) and 1(b).
+Table 3 shows the calculated f-scores for the VLMs, custom classifier, and baseline models per class and as an average over all classes.
+Appendix J shows the confusion matrices of the custom classifiers and VLMs.
+All metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in the dataset.
+Table 3: F1-scores per-class for zero-shot VLMs (preferred when no training data available), 3-fold cross validated f-scores for the custom architecture (when training data and moderate computation is available), and 5-fold cross validated f-scores for baseline classifiers (when training data and limited computation is available).
+zero-shot VLMs Classifier Baseline Classifiers
+Class LLaVA GPT-4o smolVLM Qwen DT GNB KNN SVM
+Evacuees 17.65 37.21 25.84 12.70 88.55±3.20 49.42±6.66 44.44±4.65 44.39±4.50 38.42±3.18
+General Information 19.05 20.37 00.00 21.88 82.04±3.54 32.49±4.63 38.05±8.41 32.88±6.67 39.35±10.31
+Preparedness 57.14 62.50 22.58 67.69 66.02±4.10 48.04±6.96 62.55±5.53 55.69±5.91 54.51±2.33
+Weather Reports 64.52 62.39 00.00 50.00 89.01±4.51 45.27±6.21 61.70±4.80 64.31±5.07 66.69±4.24
+Warnings & Status Updates 34.94 48.00 00.00 42.73 93.19±2.02 43.35±1.27 45.87±3.35 59.32±2.68 56.99±2.14
+Reports of Actions of Responders 26.09 69.09 00.00 63.89 81.10±2.14 49.43±2.23 51.37±4.50 54.35±6.08 61.72±5.73
+Infrastructure 36.36 37.50 00.00 50.00 87.88±1.86 51.62±5.27 66.32±5.60 68.85±4.35 70.37±3.81
+Political 69.09 77.86 00.00 85.25 83.73±0.29 51.58±5.78 66.59±4.43 55.02±5.48 60.53±4.27
+Insurance 88.52 86.15 93.75 90.91 87.01±1.22 87.25±5.10 90.96±4.03 87.27±5.24 92.72±3.97
+Advertisement 08.33 73.17 00.00 45.16 89.37±1.71 49.42±6.66 47.84±5.22 64.63±6.17 63.96±11.29
+Smoke & Air Quality 72.46 75.38 45.48 65.96 84.20±6.11 57.79±3.33 71.38±1.20 74.93±1.17 76.24±2.51
+Support 27.27 56.52 05.86 43.84 77.72±2.26 38.76±7.23 59.99±7.05 45.73±6.62 44.34±9.21
+Other 05.22 57.58 25.36 52.46 81.29±1.00 40.44±4.48 53.68±5.21 50.58±2.45 56.33±2.27
+Weighted Average 44.69 61.17 19.86 55.94 84.48±0.69 49.30±2.01 59.56±1.10 60.90±0.80 62.65±1.29
+The custom classifier achieved the highest weighted average F1-score (84.48%), outperforming both zero-shot vision-language models and traditional baseline classifiers across all classes.
+From the results in Table 3 and Appendix J, the ‘Evacuees’ class had the lowest F1-score under the custom classifier (78.55%), likely due to high variability in both imagery and associated text.
+On the other hand, the ‘Insurance’ class achieved the highest performance (91.86%), reflecting the strong visual and textual consistency across samples.
+Other high-performing classes include ‘Political’ (89.50 ‘Smoke& Air Quality’ (89.35%), and ‘Advertisement’ (89.17%), which may be more visually or textually distinct.
+In contrast, mid-range scores are seen in classes like ‘Support’ (77.12%) and ‘Warnings & Status Updates’ (79.31%), which often contain dense textual overlays that may not align well with the image modality.
+6 Discussion
+Task-specific fine-tuning is critical: Table 3 shows that classical models fine-tuned on labeled data generally outperform zero-shot VLMs, with the custom model achieving  23% higher f-score than VLMs.
+Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix G.
+VLMs involve tradeoffs: Zero-shot VLMs, while being a practical option when no labeled data is available, pose challenges in accuracy and cost.GPT-4o-mini offers strong results, but is costly and lacks transparency.
+Open-source models like LLaVA, Qwen, and smolVLM are more accessible but generally underperform, though Qwen is a notable exception, and trails GPT-4 by just 5%.
+Overall, VLMs may offer a solution when labeled data is unavailable, but their performance is comparable to traditional models such as SVM and KNN, which require training data but can be implemented with minimum computation.
+Future Work: The dataset exhibits a significant class imbalance, and future work may involve steps to address this issue.
+Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research Sirbu et al.
+(2022); Sîrbu et al.
+(2025).
+Fine-tuning of VLMs is also a promising avenue which should be explored in future works Duan et al.
+(2024); Zhai et al.
+(2024).
+7 Analysis on Trends in Data
+Here, we present an analysis of an additional 46,279 posts from X, posted between 2018 and 2024, using the custom classifier described in Section 4.
+The collection methodology was the same as explained in section 3.1, where tweets were gathered using a curated set of wildfire-related hashtags specific to each year, selected based on known wildfire events and hotspot locations across affected regions.
+This was combined with a keyword-based search, and then the data was combined, and duplicates were removed.
+The full set of search queries can be seen in Appendix K.
 Figure 2: Distribution of collected tweets based on author location.
-Using our best-performing classifier with an F1-score of 83%, we obtained predictions for the new data.
+Figure 3: Weekly post distributions: (left) overall yearly trend, (right) class-wise distribution in 2023.
+The author locations were extracted from the samples with the goal of obtaining the province-level region from which each sample was posted.
+These could be either automatically generated or user-provided.
+The queries were run through OSMNX [^5] https://osmnx.readthedocs.io/en/stable/, which allowed for the automated search of cities or towns that users provided.
+If a user’s location was not found, or not in Canada, it was labeled as such.
+The distribution of posts by province is shown in Figure 2, indicating that the majority of the data originates from British Columbia and Alberta.
+Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix L.
+Using the custom classifier, predictions were obtained for the new data.
 This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample.
-This of course, will not uphold the same gold standard of labeling that a human annotator could achieve, and manual labeling will have to be done in the future.
-Individual classifications should be approached with skepticism, but this method should still allow for general trends in the data to be analyzed on a broader scale.
+This of course, will not uphold the same gold standard of labeling that a human annotator could achieve.However, while individual classifications may not always be accurate, general trends in the data can be analyzed and linked to real events.
 Figure 3 (left) shows the total number of posts per week sorted by year.
@@
 Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
-Figure 3: Weekly post distributions: (left) overall yearly trend, (right) class-wise distribution in 2023.
-7 Discussion
-This work benchmarks VLMs on a new dataset.
-A critical consideration in evaluation of the VLMs is that WildFireCan-MMD establishes a novel benchmark for evaluating these models.
-Most evaluation research works are performed with pre-existing datasets, and these datasets are also highly likely to be used for training VLMs and do not represent a genuine test set.
-Our newly created dataset evaluates the real generalization capabilities of such models since it was not available to the creators of these models during training.
-This is important for information management during crises because new disasters have unique requirements.
-The use of VLMs involves tradeoffs.
-Practical implications are involved in benchmarking with VLMs in a zero-shot setting.
-We discuss these implications from three perspectives: accuracy, cost and compliance with the responsible AI principles.
-The closed-source model, gpt-4o-mini, while being accurate to some extent, is expensive for large-scale deployment.
-Also, closed-source models often fail to comply with responsible AI principles due to a lack of transparency, privacy, and limited scrutiny and ethical oversight Hua et al.
-(2024).
-On the other hand, despite the free and open-source nature of LLaVA, its performance falls short in critical tasks such as disaster management.
-Also, smolVLM, while being both open-source and computationally affordable, has a performance far below acceptable rates.
-Therefore, when labeled data is unavailable, VLMs might be considered as a zero-shot classifier, but in practice, these models are not particularly well suited to this specific type of task, emphasizing the necessity for a custom fine-tuned model.
-Fine-tuning on task-specific data is essential.
-Results presented in Table 2 demonstrate that even simple approaches such as decision tree and Gaussian NB, when trained with task-specific labeled data, provide similar or better results than what is achieved from zero-shot VLMs.
-Even though our dataset is modest in size, it clearly showcases the benefits of developing a labeled dataset and fine-tuning a custom classifier.
-Further, our results show that high degrees of accuracy can be achieved for this task by fine-tuning common classification structures, such as pre-trained transformer-based text and image encoders.
-These findings align with other works such as Nadeem et al.
-(2025).
-Custom classifiers can be used to discover social media use patterns.
-The primary function of the custom classifier is to perform real-time labeling of social media data during wildfire events, enabling the timely extraction of actionable information to support disaster response efforts.
-Beyond real-time deployment, our post-hoc analysis of a large unlabeled dataset illustrates the classifier’s capacity to identify latent patterns and temporal trends, providing a scalable tool for retrospective analysis and evidence-based policy development.
-To enhance the accuracy and robustness of our model, future work can consider the application of different fusion strategies, such as multiplicative fusion, which has been shown to better capture the relationship between different modalities Koshy and Elango (2023).
-Fine-tuning of VLMs is also a promising avenue of improving the task which should be explored in future works Duan et al.
-(2024); Zhai et al.
-(2024).
 8 Conclusion
-Despite the advancements in general-purpose VLMs, they struggle to perform reliably in high-stakes, context-rich scenarios like disaster response.
-This low accuracy, alongside the resource-intensive nature of such models, necessitates developing smaller fine-tuned models, specialized for specific regions and contexts.
-To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation.
-Unlike prior disaster datasets, ours is tailored exclusively to wildfire contexts, making it a valuable benchmark for evaluating true generalization in VLMs, and suitable for training specialized models.
-As climate-induced wildfires grow in frequency and complexity, this dataset supports the development of robust, transparent, and adaptive AI tools for detection, coordination, and recovery.
-9 Acknowledgment
+Despite the existence of other natural disaster datasets from social media, these datasets often lack wildfire data from a significant period of time, typically spanning only a single month surrounding a particular event.
+Furthermore, even less of this data comes from Canada and is mainly restricted to the US context.
+To address this gap, we present WildFireCan-MMD, a wildfire-specific, multimodal dataset of social media posts curated from X, accompanied by a domain-focused taxonomy derived through multimodal topic modeling and manual annotation by multiple independent annotators.
+Unlike existing disaster datasets, WildFireCan-MMD is specifically designed for the Canadian wildfire context and provides a rich resource for training and evaluating specialized models.
+It also serves as a true test set for evaluating VLMs in a zero-shot setting.
+We provide the results of classifiers trained on this dataset from classical approaches to custom deep learning architectures and several VLMs, to showcase various scenarios in terms of data and computational resource availability.
+Our best results are obtained with an architecture that utilizes pre-trained encoders and transformer fusion for classification on this dataset, which highlights the importance of a labelled dataset.
+As climate-induced wildfires grow in frequency and complexity, this dataset and model support the development of robust tools for detection, coordination, and recovery.
+Acknowledgment
 This work was funded by the National Program Office of the National Research Council Canada, under the New Beginning Initiative Program.
 Limitations
-While our study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
+While this study introduces a novel wildfire-specific multimodal dataset and demonstrates the effectiveness of fine-tuned task-specific models, several limitations should be acknowledged.
 The first limitation is the scope and size of the dataset.
 Although WildFireCan-MMD provides a targeted and realistic benchmark for wildfire-related content, its size remains modest compared to large-scale datasets used to train or evaluate deep learning models.
-This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of our results to other wildfire events or regions.
+This limits the diversity of linguistic and visual expressions captured and may affect the generalizability of these results to other wildfire events or regions.
 Platform and geographic bias might be present in our data.
@@
 Deploying these models in real-world disaster contexts without appropriate oversight, interpretability mechanisms, or human-in-the-loop verification may lead to misclassification, misinformation amplification, or missed signals, potentially affecting emergency coordination and public trust.
-Appendix A Data collection Queries
+Appendix A Data Collection
 Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
@@
 Table A.1: The queries used for collecting the data for WildFireCan-MMD.
-Appendix B Topic modeling parameters
-In order to optimize the performance of BERTopic, careful consideration of hyperparameters is key McInnes et al.
-(2018).
-First, text and image inputs are vectorized using the pre-trained “Clip-ViT-B-32” [^11] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model.
-Then, BERTopic uses the UMAP algorithm for dimension reduction, as it keeps a significant portion of the high-dimensional local structure in lower dimensionality.
-We use the default parameter settings of BERTopic, which reduces the dimensionality to 5, while keeping the size of the local neighborhood at 15.
-After dimension reduction, BERTopic uses HDBSCAN to cluster the vectors.
-HDBSCAN is a density-based algorithm that does not force data points to clusters, as it can consider them outliers.
-HDBSCAN has been shown to be effective when paired with UMAP, since UMAP maintains the local structure in the lower-dimensional space.
-After forming the clusters, BERTopic extracts keywords from each topic to uncover their central theme.
-For that, we set the parameters so that stop words are removed and both unigrams and bigrams are considered as candidates of topic keywords.
-Once we had reduced our topics to around 15, it became clearer which topics were the strongest in our dataset, and we were able to begin to see the humanitarian utility of some of these topics.
-Some examples of useful topics were posts about smoke and air quality, weather reports, and reports of actions of first responders.
-We also noticed some less useful posts, such as advertisements and political debates/complaints.
-Some examples of these topics, as well as their corresponding keywords, can be seen in Figure B.1.
+Appendix B Annotation
+Figure B.1 shows the distribution of the labels over the classes by each independent annotator.
+Figure B.1: Counts of labels by each annotator
+Appendix C Topic Modeling
+Figure C.1 shows examples of topics extracted by the unsupervised BERTopic.
+We present representative images and corresponding keywords generated by BERTopic for each example.
+As seen in the captions of the image, some of the topics are not useful and present non-coherent concepts, while others are useful and relevant to wildfires.
 Keywords: [crews, safe, poles, damaged, continue, lines, working restore, power, conditions, restore]
@@
 (c) Non-useful
-Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
-Appendix C Prompts for zero-shot classification
-We show the full prompt used to evaluate the aforementioned VLMs in Table C.1. The following hyperparameters are used for all VLMs: temperature=0.1, max_new_tokens=1024.
-Table C.1: System and User prompt used for VLMs.
+Figure C.1: Examples of useful/non-useful topics from topic modeling on our dataset.
+Appendix D Prompts for Zero-Shot Classification
+The full prompt used to evaluate the aforementioned VLMs is shown in Table D.1.
+Table D.1: System and User prompt used for VLMs.
 System Prompt:
@@
 You may only answer with the chosen option’s letter."
-Appendix D Structures of custom classifiers
-Figures D.1 and D.2 show the structures of the finetuned models. We described these models in detail in Section 4.3.
-Figure D.1: Custom multimodal 2-head early fusion classifier structure.
-Figure D.2: Custom multimodal 3-head early fusion classifier structure.
-Appendix E Confusion matrices of classifiers
-For deeper insights into classifiers’ performances, we show the confusion matrices of all classifiers in Figures E.1 and E.2.
-(a) 2-head classifier.
-(b) 3-head classifier.
-Figure E.1: Confusion matrices of custom classifiers.
+Appendix E Baseline Classifiers
+To establish a baseline, several classical methods were evaluated on the dataset, namely DT, GNB, KNN, SVM.
+For encoding data into numerical features for these methods, (pretrained, but not finetuned) RoBERTa and ViT were used to extract feature sets for each modality.
+The concatenated feature sets from the last hidden state were saved.
+Principle Component Analysis (PCA) was used on the feature sets to reduce the dimensionality from 768 to 250 for the SVM method as it improved the results.
+For DT, class_weight=’balanced’ was used.
+For KNN, n_neighbors=1 was used.
+All other hyperparameters remained default from Scikit-learn [^6] https://scikit-learn.org/stable/.
+Five-fold cross validation was used to train and test all methods to report a confidence interval.
+Appendix F VLMs as Zero-Shot Classifiers:
+Four VLMs were considered for experimentation, GPT-4o-mini Achiam et al.
+(2023), LLaVA Liu et al.
+(2023), smolVLM Marafioti et al.
+(2025), Qwen2.5-VL Bai et al.
+(2025) and tested them in a zero-shot setting.
+Classification prompts for all three VLMs were created.
+With each query to the model, a list of all categories and their descriptions are given as a prompt along with a data instance (text + image) to classify.
+The prompt specifies that the model’s task is to choose one of the provided categories and respond only with its corresponding letter.
+The system and user prompts can be seen in Appendix D and are the same for all VLMs.
+The same hyperparameters are used for all VLMs with temperature=0.1 num_beams=1, max_new_tokens=1024.
+The same test set was used for all VLMs and consisted of a stratified split with random_state=8.
+This single split was discovered to be a good split from previous experimentation with baseline and custom classifiers, and was necessary due to the cost of running the VLM experiments multiple times.
+SmolVLM-2B: Being the smallest of the VLMs experimented with, with only 2.2B parameters and <10GB of GPU ram, experiments were able to be performed locally without quantization.
+Qwen2.5-VL-7B: Qwen is an intermediately-size VLM with around 7B parameters.
+Being too large for the hardware available, manual quantization using BitsandBytes was performed to shrink the precision down to 4-bit to be used locally.
+LLaVA-13B: Also considered an intermediately-sized option, the 13B 4-bit quantized version of LLaVA was used due to hardware constraints.
+Using a combination of a publicly available Colab notebook  [^7] https://github.com/camenduru/LLaVA-colab, the author of which performed the quantization, and the original code from the open-source LLaVA repository  [^8] https://github.com/haotian-liu/LLaVA, the model was able to be used with Google Colab.
+GPT-4o-mini: The largest of the VLMs tested, a paid API service had to be used for querying GPT-4o-mini 999version:GPT-4o-mini-2024-07-18, and it is also the only closed-source model tested.
+A specified content field in each query contains the system prompt, text and image from the tweet, and user prompt.
+For each method, the responses are saved as a file where each line contains the model’s answer supplied as a letter from the options shown in Table D.1.
+This then allows for the calculation of metrics.
+Appendix G Related Architectures
+Table G.1 briefly describes the architectures proposed by various authors on the CrisisMMD dataset.
+The purpose of this table is to demonstrate the literature review used to select the classifier proposed in section 4 and give an idea of the expected accuracy of such a classifier on this task of wildfire social-media data.
+Paper Method Task 1 Task 2 Task 3
+Agar. 2020 Agarwal et al. (2020) InceptionV3+RCNN+AttentionFusion 99.00 99.00 97.00
+Abav. 2020 Abavisani et al. (2020) DenseNet+BERT+CrossAttentionFusion 89.35 91.82 70.41
+Ofli. 2020 Ofli et al. (2020) VGG16+CNN+Concat.Fusion 84.20 78.30 -
+Xukun 2020 Li and Caragea (2020) AlexNet+GRU+Concat.Fusion 78.13 - -
+Madi. 2021 Madichetty et al. (2021) DenseNet+BERT+AverageFusion 74.85 - -
+Kumar 2022 Kumar et al. (2022) VGG16+LSTM+Concat.Fusion 82.00 - -
+Pran. 2022 Pranesh (2022) ResNet+RoBERTa+TransformerAttentionFusion 85.10 89.50 -
+Sirbu 2022 Sirbu et al. (2022) semi-supervised 82.00 87.20 88.70
+Diva. 2022 Divakaran et al. (2022) DenseNet+RoBERTa+Concat.Fusion 89.26 85.87 -
+Kotha 2022 Kotha et al. (2022) RegNetY320+BERT+Concat.Fusion 89.63 86.59 -
+Khat. 2022 Khattar and Quadri (2022) VGG16+BiLSTM+CrossAttentionFusion 84.08 - -
+Madi. 2023 Madichetty et al. (2023) VGG16+RoBERTa+MultiplicativeFusion 90.21 - -
+Basit 2023 Basit et al. (2023) DeIT+ALBERT+Concat.Fusion 79.00 80.00 86.00
+Koshy 2023 Koshy and Elango (2023) ViT+RoBERTa+BiLSTM+BilinearFusion 97.14 - -
+Gupta 2024 Gupta et al. (2024) DenseNet+ELECTRA+CrossAttentionFusion 91.20 93.40 72.20
+Shet. 2025 Shetty et al. (2025) ViT+RoBERTa+CrossAttentionFusion 91.25 89.63 68.72
+Sirbu 2025 Sîrbu et al. (2025) semi-supervised 91.22 89.55 -
+Teng 2025 Teng and Öhman (2025) VGG16+BERT+CrossAttentionFusion 81.00 - -
+Dar 2025 Dar et al. (2025) (ResNet50+BERT)+(Graph)+IDEAFusion 98.23 90.13 -
+Table G.1: Literature review of deep-learning approaches used by other works on the CrisisMMD dataset.
+Appendix H Ablation & Other Experiments
+Tables H.1 and H.2 show the different text/image encoders experimented with, using either only text or only images respectively.
+All experiments here were conducted using the same train/test split, which was earlier identified as being a good split through cross validation experiments on different random seeds.
+Experiments were done on different final feature representations by using either mean pooling of the last hidden state or, using the CLS token.
+Also, freezing the first half of the layers as this has been shown to increase accuracy Ingle et al.
+(2022).
+All encoders had a linear layer appended for classification.
+For text, "roberta-base" performed best when using just the CLS token for classification and freezing the first half of the encoder layers, as shown in Table H.1.
+Later experimentation found that the unfrozen version performed better in the final classifier.
+For images, "vit-base-patch16-384" performed best when using just the CLS token for classification without any freezing.
+Many of the fusion experiments tested, as well as some architectures from other studies can be seen in Table H.3.
+The same train/test split was used as before for unimodal classification.
+The best result, belonging to the selected classifier, is shown in bold.
+Table H.4 shows the ablation study performed on the selected model.
+(1) is the result of the selected model, (2) is the result of the model with a simpler classification head consisting of a single linear layer, (3) is the model with the proposed classification head but without the fusion module,
+Three-fold cross-validation was performed later to support the result as described in section 4.
+Text Model Notes F1-Weighted
+’bert-base-cased’ mean-pooling 0.7779
+’bert-base-cased’ CLS 0.7793
+"bertweet-base" mean-pooling 0.7900
+"bertweet-base" CLS 0.7927
+"ModernBERT-base" mean-pooling 0.7934
+"ModernBERT-base" CLS 0.7862
+"roberta-base" robertaforseq 0.7718
+"roberta-base" mean-pooling 0.7934
+"roberta-base" mean-pooling, half-frozen 0.7940
+"roberta-base" mean-pooling, custom vocab 0.7209
+"roberta-base" CLS 0.7798
+"roberta-base" CLS, half-frozen 0.8008
+Table H.1: Text-only results.
+Image Model Notes F1-Weighted
+DenseNet121 ImageNet weights 0.5733
+InceptionV3 ImageNet weights 0.5804
+ResNet50 ImageNet weights 0.5841
+VGG16 ImageNet weights 0.5993
+"facebook/deit-base-patch16-384" CLS 0.6260
+"google/vit-base-patch16-384" half-frozen, CLS 0.6290
+"google/vit-base-patch16-384" mean-pooling 0.6404
+"google/vit-base-patch16-384" CLS 0.6419
+Table H.2: Image-only results.
+Text Model Image Model Fusion Type Notes F1
+RoBERTa ViT Cross Attention encoders pretrained individually, q=image k=text v=text 0.7855
+RoBERTa ViT Cross Attention encoders pretrained individually, q=text k=image v=image 0.6489
+RoBERTa Coca Concatenative COCA/RoBERTa individually aligned w/ description of the label 0.1587
+RoBERTa+BiLSTM ViT BiLSTM+BiLinear Koshy’s model Koshy and Elango (2023) 0.6149
+CLIP CLIP "openai/clip-vit-base-patch32" 0.7072
+VILT VILT "dandelin/vilt-b32-finetuned-nlvr2" 0.7638
+RoBERTa ViT Concatenative ViT/RoBERTa individually aligned w/ description of the label 0.6507
+RoBERTa ViT Transformer transformer fusion: h_dim 1024 and 0.1 dropout 0.8263
+RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout 0.8364
+RoBERTa ViT Transformer RoBERTa/ViT fully frozen 0.6164
+RoBERTa ViT Concatenative stopwords removed from text before RoBERTa 0.7870
+RoBERTa ViT Concatenative CLS into single classification layer 0.8249
+RoBERTa x2 ViT Concatenative extract text from images with OCR 0.8102
+ModernBERT ViT Transformer transformer fusion: hidden dim 2048 and 0.2 dropout 0.8149
+ModernBERT ViT Concatenative CLS into single classification layer 0.7966
+RoBERTa ViT Concatenative mean-pooling of last hidden states into linear classification layer 0.7911
+RoBERTa ViT Concatenative roberta:mean, vit:cls 0.8028
+RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout RoBERTa:mean, ViT:cls 0.8223
+RoBERTa ViT Cross Attention same as proposed but with cross attention instead of transformer 0.8208
+Table H.3: Multimodal results.
+Ablation Explanation F1-Weighted
+1 Proposed 83.64
+2 Linear class. head 82.70
+3 No fusion module 81.71
+Table H.4: Ablation of model components.
+Appendix I Structure of Custom Classifier
+Figure I.1 shows the architecture of the selected approach. This model is described in detail in Section 4.
+Figure I.1: Selected dual-encoder architecture utilizing transformer fusion module.
+Appendix J Confusion Matrices
+For deeper insights into the classifiers performances, the confusion matrices are shown in Figures J.1 and J.2.
+(a) Random seed: 8
+(b) Random seed: 12
+(c) Random seed: 14
+Figure J.1: Confusion matricies for selected classifier over the different train/test splits.
 (a) LLaVA.
@@
 (c) smolVLM.
-Figure E.2: Confusion matrices of VLMs.
-Appendix F Queries Used for Additional Data Collection :
-Below is the list of queries and keywords used to collect additional unlabeled data discussed in Section 6.
+(d) Qwen.
+Figure J.2: Confusion matrices of VLMs.
+Appendix K Queries Used for Additional Data Collection
+Shown here is the list of queries and keywords used to collect the additional unlabeled data discussed in Section 7.
 2018
@@
 Keyword Query: (canada OR ontario OR quebec OR sasksatchewan OR manitoba OR northwest territories OR yukon OR british columbia OR alberta OR jasper) (wildfire OR forest fire)
-Note: Every query above includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
-Appendix G Unlabeled dataset
-A word cloud of the text data of the unlabeled dataset collected in Section 6, and also the frequency distribution of the posts by word count, are shown in Figures 1(a) and 1(b).
+Note: Every query includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
+Appendix L Unlabeled Dataset
+Figure 2 shows the distribution of post origins.
+A word cloud of the text data of the unlabeled dataset collected in Section 7, and also the frequency distribution of the posts by word count, are shown in Figure L.1.
 (a) Word cloud of the posts.
 (b) Frquency distribution of the posts w.r.t their lengths.
-Figure G.1: Word cloud of the textual part of the unlabeled tweets and the frequency distribution of them regarding their word counts.
+Figure L.1: (a) Word cloud of the text from the unlabeled tweets and (b) the frequency distribution of their word counts.

### v4

@@
 In this work, we focus on multimodal wildfire social media data, which, although existing in current datasets, is currently underrepresented in Canadian contexts.
-We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes.
+We present WildFireCan-MMD, a new multimodal dataset of posts on X from recent Canadian wildfires, annotated across twelve key themes.
 We evaluate zero-shot vision-language models on this dataset and compare their results with those of custom-trained and baseline classifiers.
@@
 WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada
-Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,MarziehAmini@cunet.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca
+Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca,Efstratiosaivaliotis@cmail.carleton.ca, KhaledMslmani@cmail.carleton.ca, MarziehAmini@cunet.carleton.ca
 1 Introduction
-Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al.
-(2024b), mental health monitoring Ghanadian et al.
-(2024); Dalal et al.
-(2024), detecting and combating hate speech in digital environments Kiritchenko et al.
-(2021); Kumar et al.
-(2024a), pandemic discourse analysis Imran et al.
-(2022) and sentiment and behavioral trend analysis Ma et al.
-(2024); Curto et al.
-(2024).
-Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam et al.
-(2018); Ofli et al.
-(2020); Alam et al.
-(2021); Dwarakanath et al.
-(2021); Algiriyage et al.
-(2022); Hossain et al.
-(2022); Khattar and Quadri (2022); Basit et al.
-(2023); Koshy and Elango (2023); Rezk et al.
-(2023).
+Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi2023; kumar2024extracting, mental health monitoring ghanadian2024socially; dalal2024review, detecting and combating hate speech in digital environments kiritchenko2021confronting; kumar2024detecting, pandemic discourse analysis tbcov19 and sentiment and behavioral trend analysis ma2024toward; curto2024crime.
+Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam2018; Ofli2020; alam2021humaid; Dwarakanath2021; Algiriyage2022; Hossain2022; Khattar2022; Basit2023; Koshy2023; Rezk2023.
 Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information.
-The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston et al.
-(2015).
+The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston2015.
 This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information.
 Therefore, there is a critical need to develop an automated method for sifting through this information.
-To foster research in this area, Alam et al.
-(2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes.
+To foster research in this area, Alam2018; alam2021humaid developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes.
 However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios.
@@
 Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising 1589 samples, and is specific to the California wildfires of October 2017.
-The wildfire-related samples from Canada in the HumAID dataset Alam et al.
-(2021) are limited to a single month in 2016, as well as another month of California samples from 2018.
+The wildfire-related samples from Canada in the HumAID dataset alam2021humaid are limited to a single month in 2016, as well as another month of California samples from 2018.
 Therefore, wildfires in Canada remain underexplored in social media analysis research.
@@
 To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several Vision-Language Models (VLMs), custom deep-learning classifiers, as well as classical approaches, as a minimum achievable performance.
-Our best-performing classifier utilizes RoBERTa Liu (2019) and Vision Transformer (ViT) models to jointly process the text and image components of tweets.
+Our best-performing classifier utilizes RoBERTa roberta2019 and Vision Transformer (ViT) models to jointly process the text and image components of tweets.
 A custom transformer-based fusion module is designed to enable the modalities to interact deeply.
@@
 Using our best-performing classifier, we analyze this unlabeled data and uncover trends and patterns of social media use during wildfires in Canada.
-The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at [Anonymized GitHub], under CC-BY License.
+The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available on Github (https://github.com/Multimodal-Social-Media-Data-Analysis/WildfireCanMMD).
 The contributions of this study are summarized as:
@@
 Significant research has been conducted on the role of social media during disasters.
-Fauzi (2023) conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions.
-Acikara et al.
-(2023) highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information.
-Muniz-Rodriguez et al.
-(2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster.
-In other works, Dwarakanath et al.
-(2021), Bukar et al.
-(2020), and Mundottukandi et al.
-(2024) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities.
-Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage et al.
-(2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
-Our work is motivated by this literature as well as the literature shown in Appendix G which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
+Fauzi2023 conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions.
+Acikara2023 highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information.
+Muniz2020 focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster.
+In other works, Dwarakanath2021, 9220091, and 10485423 discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities.
+Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage2022 discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
+Our work is motivated by this literature as well as the literature shown in Appendix F which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
 2.2 Multimodal classification for disaster response
 We review this literature based on the available datasets and previously explored algorithms.
-Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters Olteanu et al.
-(2014); Dong et al.
-(2021); Ochoa and Comes (2021), it is now generally considered that multimodal approaches offer stronger results Ofli et al.
-(2020); Abavisani et al.
-(2020); Ramirez et al.
-(2025).
-The study by Alam et al.
-(2018) released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes).
+Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters crisislex2014; Dong2021; Ochoa2021, it is now generally considered that multimodal approaches offer stronger results Ofli2020; Abavisani2020; ramirez2025.
+The study by Alam2018 released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes).
 CrisisMMD contains thousands of tweets labeled for categories like “infrastructure and utility damage” or “rescue volunteering or donation effort”.
-Additionally, Alam et al.
-(2018) later released the HumAID Alam et al.
-(2021) dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019).
+Additionally, alam2021humaid later released the HumAID dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019).
 It is similarly annotated for emergency response categories (e.g.
 casualties, infrastructure damage).
-There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex Olteanu et al.
-(2015) and CrisisMMD Alam et al.
-(2018) datasets Wu et al.
-(2022).
+There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex crisislext262015 and CrisisMMD datasets Xuehua2022.
 CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images.
-It remains the most prominent dataset in this domain, used in many recent studies Ofli et al.
-(2020); Khattar and Quadri (2022); Basit et al.
-(2023); Koshy and Elango (2023); Rezk et al.
-(2023) and many others.
+It remains the most prominent dataset in this domain, used in many recent studies Ofli2020; Khattar2022; Basit2023; Koshy2023; Rezk2023 and many others.
 Therefore, there is a notable gap in the in the availability of multimodal wildfire datasets.
@@
 Additionally, this data is only from the California wildfires of October 2017.
-The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016, as well as another month of California samples from 2018.
+The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016 (2,259 samples), as well as another month of California samples from 2018 (7,444 samples).
 Therefore, wildfires in Canada remain underexplored in social media analysis research.
 To address this limitation, this work focuses on gathering social media data related to Canadian wildfires, and proposes WildFireCan-MMD, a new Canadian wildfire social-media dataset with a wildfire-specific taxonomy 12 categories.
-Algorithms: Previous work by Ofli et al.
-(2020) introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality.
+Algorithms: Previous work by Ofli2020 introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality.
 Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures respectively, followed by a dense layer for prediction.
-The work by Basit et al.
-(2023), expands on Ofli et al.
-(2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work.
-They found that using a combination of AlBERT Lan et al.
-(2020) as a text embedder, and DeIT Touvron et al.
-(2021) as an image embedder, provided the best results for the eight-class classification task.
-Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs.
+The work by Basit2023, expands on Ofli2020 by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work.
+They found that using a combination of AlBERT albert2020 as a text embedder, and DeIT deit2021 as an image embedder, provided the best results for the eight-class classification task.
+Koshy2023 implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs.
 Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods.
-Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al.
-(2018).
-Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix G.
+Specifically, they demonstrated that using the fine-tuned RoBERTa roberta2019 model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam2018.
+Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix F.
 Most previous studies use a CNN for the image modality and then a BERT-based transformer for the text modality.
@@
 Cross-attention fusion of the two modalities seems to be another popular approach that seems to give promising results.
-The approach given by Pranesh (2022) proposes a transformer-like structure to fuse the two modalities.
+The approach given by pranesh2022 proposes a transformer-like structure to fuse the two modalities.
 We build on the this previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
-In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al.
-(2024); Li et al.
-(2022).
-Specifically, we evaluate three VLMs, GPT-4 Achiam et al.
-(2023), LLaVA Liu et al.
-(2023), and SmolVLM Marafioti et al.
-(2025), to assess performance across a range of capabilities and computational requirements.
-GPT-4 serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
+In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Duzhen2024; Li2022.
+Specifically, we evaluate four VLMs, GPT-4o-mini OpenAI2023, LLaVA-v1.5-13B Liu2023, Qwen2.5-VL-7B qwen2025, and SmolVLM-2B smolvlm2025, to assess performance across a range of capabilities and computational requirements.
+GPT-4o-mini serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
 3 The WildFireCan-MMD Dataset
@@
 The dataset comprises 4,688 unique image-text pairs gathered from the wildfire season (April-September) of each respective year.
-Data collection was conducted using a hashtag-based search strategy as seen in other works Alam et al.
-(2018); Gupta and Hewett (2020); Cruickshank and Carley (2020).
-Appendix A presents the full queries used for data collection.
+Data collection was conducted using a hashtag-based search strategy as seen in other works Alam2018; gupta2020; Cruickshank2020.
+Table 1 presents the full queries used for data collection.
 To curate the hashtags, basic hashtags such as "#wildfire, #forestfire" were first used.
@@
 The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region from which the tweet originated.
-3.2 Taxonomy Development
-To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst (2022).
-Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic.
-The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15.
-Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords.
-Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics.
-Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased.
-For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected.
-Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear.
-Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders.
-Less useful posts were also noticed, such as advertisements and political debates/complaints.
-Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix C.
-The topics were inspected and 12 classes were decided upon.
-The full list of classes and their descriptions is given in Table 1.
-To ensure the reliability of the labeling process, the definitions and boundaries of each category were discussed among the research team and a shared understanding was reached.
-Through this discussion, it was determined that the classes are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image).
-This discussion, along with unsupervised topic modeling (BERTopic) were used to find strong topics which were then later assigned to a class.
-Also, the selection of classes was guided by the topics created by BERTopic, observations made through manual annotation, and the perceived relevance of classes to stakeholders in disaster scenarios.
-For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel.
-The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality.
-The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
-An ‘Other’ class is added to accommodate the samples that do not fit into any of the 12 classes.
+Year Query
+2022/2023 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
+2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
+Table 1: The queries used for collecting the data for WildFireCan-MMD.
 Table 2: Categories of WildFireCan-MMD, their descriptions and total count of each class.
@@
 Other No ’useful’ information, focusing on images (e.g., scenery), general complaining, or irrelevant content. 507
-3.3 Data Labeling
-Labeling was carried out in-house by three independent annotators using Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 1.
-Examples of labeled tweets are presented in Figure 1.
-Each image-text pair was given a single label.
-In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample.
-For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’.
-In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone.
-If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided.
-The size of labeled classes is presented in Table 1.
-Annotation was done by three independent annotators, all authors of this paper.
-The annotators are male, 20-30 year old, fluent in English and reside in Canada.
-First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes.
-Over the course of labeling, samples were flagged that contained personal information.
-A vote between the labels from the three annotators decided the final labels.
-Annotation agreement scores can be seen in Table 2.
-A majority agreement was reached on 88.4% of the labels.
-Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels.
-From the results, annotators 1 & 2 had the best agreement scores.
-This is further demonstrated by Cohen’s and Fleiss’ Kappa metrics, which indicate that a moderate-to-good level of agreement is achieved between the independent annotators.
-In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators.
-Additional statistics can be found in Appendix B.
+3.2 Taxonomy Development
+To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst2022.
+Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic.
+The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15.
+Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords.
+Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics.
+Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased.
+For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected.
+Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear.
+Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders.
+Less useful posts were also noticed, such as advertisements and political debates/complaints.
+Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix B.
+The topics were inspected and 12 classes were decided upon.
+The full list of classes and their descriptions is given in Table 2.
+To ensure the reliability of the labeling process, the definitions and boundaries of each category were discussed among the research team and a shared understanding was reached.
+Through this discussion, it was determined that the classes are well-defined and grounded in observable cues (e.g., specific topics, hazards, or references in the text or image).
+This discussion, along with unsupervised topic modeling (BERTopic) were used to find strong topics which were then later assigned to a class.
+Also, the selection of classes was guided by the topics created by BERTopic, observations made through manual annotation, and the perceived relevance of classes to stakeholders in disaster scenarios.
+For example, ‘Warnings & Status Updates’, ‘Reports of Actions of Responders’, and ‘Infrastructure’ could be useful to first responders and emergency personnel.
+The ‘Evacuees,’General Information,’ and ‘Preparedness’ categories could be useful to the general public to be informed about the events happening around their locality.
+The ‘Political’ category could be useful to the media to gauge public opinion and policies on events surrounding the disaster.
+An ‘Other’ class is added to accommodate the samples that do not fit into any of the 12 classes.
 Table 3: Annotation agreement statistics.
@@
 Fleiss’ Kappa 54.3%
+3.3 Data Labeling
+Annotation was performed in Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 2.
+Examples of labeled tweets are presented in Figure 1.
+Each image-text pair was given a single label.
+In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample.
+For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’.
+In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone.
+If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided.
+The size of labeled classes is presented in Table 2.
+Samples were annotated by three independent annotators, all authors of this paper.
+The annotators are male, 20-30 year old, fluent in English and reside in Canada.
+First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes.
+Over the course of labeling, samples were flagged that contained personal information.
+A vote between the labels from the three annotators decided the final labels.
+Annotation agreement scores can be seen in Table 3.
+A majority agreement was reached on 88.4% of the labels.
+Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels.
+From the results, annotators 1 & 2 had the best agreement scores.
+This is further demonstrated by Cohen’s Kappa metric between the different pairs of annotators.
+Fleiss’ Kappa metric however, indicates that a moderate level of agreement is achieved between the independent annotators.
+In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators.
+Additional statistics can be found in Appendix A.
 Label: Reports of Actions of Responders
@@
 We investigate three categories of classification methods.
-First, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers.
+First, classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) are used to investigate a case where labeled data is available but computation is limited.
+Then, several VLMs such as LLaVA, Qwen, smolVLM, and GPT-4o-mini are employed as zero-shot classifiers.
 This is the right choice when no training data is available.
-Then we examined classical methods such as Decision Trees (DT), Gaussian Naive Bayes (GNB), K-Nearest Neighbours (KNN) and Support Vector Machines (SVM) to investigate a case where labeled data is available but computation is limited.
 Finally, we experimented with a fine-tuned transformer-based classifier to assess its performance in the presence of labeled data and moderate computational resources.
-Further information on classical methods and VLMs can be found in Appendices E and F.
-Architecture of custom classifier: To choose the architecture of the custom classifier, we experimented with the architectures proposed by various authors on the CrisisMMD dataset (presented in Appendix G) and evaluated them on the WildFireCAD-MMD dataset.
+Further information on classical methods and VLMs can be found in Appendices D and E.
+Architecture of custom classifier: To determine which components to use in the selected architecture, we first experimented with various unimodal classifiers for each modality, as seen in Tables 4 & 5.
 These included CNNs and transformers for the image modality, as well as various BERT-based transformers for the text modality.
-After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in literature and included in Table G.1 were tested to improve the classification.
-More information about these experiments, as well as an ablation study, can be found in Appendix H.
+After individually determining the best model for each modality (ViT for image with f1: 64.19%, and RoBERTA for text with f1: 77.98%), various fusion techniques found in the literature and included in Table F.1 were tested to improve the classification.
+The results of these multimodal experiments can be seen in Table 6.
+More information about these experiments, as well as an ablation study, can be found in Appendix G.
 The best performing classifier consisted of two parallel encoders, an image and text encoder.
@@
 A final classification head projects the dimension from 512 to 256, which is followed by ReLU activation and dropout=0.2, then finally from 256 to num_classes.
-Figure I.1 shows the structure of this architecture.
-Training and Evaluation: Three-fold cross validation with stratified sampling was used to train the model and report the results (mean $\pm$ std).
+Figure H.1 shows the structure of this architecture.
+Training and Evaluation: Three data splits with stratified sampling were used to train the model and report the results.
 The random seeds from splitting the data were kept for reproducibility.
@@
 The model was tested on the held-out test set (938 samples).
+Table 4: Text-only classification results.
+Text Model Notes F1-Weighted
+’bert-base-cased’ mean-pooling 0.7779
+’bert-base-cased’ CLS 0.7793
+"bertweet-base" mean-pooling 0.7900
+"bertweet-base" CLS 0.7927
+"ModernBERT-base" mean-pooling 0.7934
+"ModernBERT-base" CLS 0.7862
+"roberta-base" robertaforseq 0.7718
+"roberta-base" mean-pooling 0.7934
+"roberta-base" mean-pooling, half-frozen 0.7940
+"roberta-base" mean-pooling, custom vocab 0.7209
+"roberta-base" CLS 0.7798
+"roberta-base" CLS, half-frozen 0.8008
+Table 5: Image-only classification results.
+Image Model Notes F1-Weighted
+DenseNet121 ImageNet weights 0.5733
+InceptionV3 ImageNet weights 0.5804
+ResNet50 ImageNet weights 0.5841
+VGG16 ImageNet weights 0.5993
+"facebook/deit-base-patch16-384" CLS 0.6260
+"google/vit-base-patch16-384" half-frozen, CLS 0.6290
+"google/vit-base-patch16-384" mean-pooling 0.6404
+"google/vit-base-patch16-384" CLS 0.6419
 5 Experimental Results
-Table 3 shows the calculated f-scores for the VLMs, custom classifier, and baseline models per class and as an average over all classes.
-Appendix J shows the confusion matrices of the custom classifiers and VLMs.
-All metrics reported in these tables use ’weighted’ F1-scoring to better represent the class imbalance observed in the dataset.
-Table 3: F1-scores per-class for zero-shot VLMs (preferred when no training data available), 3-fold cross validated f-scores for the custom architecture (when training data and moderate computation is available), and 5-fold cross validated f-scores for baseline classifiers (when training data and limited computation is available).
-zero-shot VLMs Classifier Baseline Classifiers
-Class LLaVA GPT-4o smolVLM Qwen DT GNB KNN SVM
-Evacuees 17.65 37.21 25.84 12.70 88.55±3.20 49.42±6.66 44.44±4.65 44.39±4.50 38.42±3.18
-General Information 19.05 20.37 00.00 21.88 82.04±3.54 32.49±4.63 38.05±8.41 32.88±6.67 39.35±10.31
-Preparedness 57.14 62.50 22.58 67.69 66.02±4.10 48.04±6.96 62.55±5.53 55.69±5.91 54.51±2.33
-Weather Reports 64.52 62.39 00.00 50.00 89.01±4.51 45.27±6.21 61.70±4.80 64.31±5.07 66.69±4.24
-Warnings & Status Updates 34.94 48.00 00.00 42.73 93.19±2.02 43.35±1.27 45.87±3.35 59.32±2.68 56.99±2.14
-Reports of Actions of Responders 26.09 69.09 00.00 63.89 81.10±2.14 49.43±2.23 51.37±4.50 54.35±6.08 61.72±5.73
-Infrastructure 36.36 37.50 00.00 50.00 87.88±1.86 51.62±5.27 66.32±5.60 68.85±4.35 70.37±3.81
-Political 69.09 77.86 00.00 85.25 83.73±0.29 51.58±5.78 66.59±4.43 55.02±5.48 60.53±4.27
-Insurance 88.52 86.15 93.75 90.91 87.01±1.22 87.25±5.10 90.96±4.03 87.27±5.24 92.72±3.97
-Advertisement 08.33 73.17 00.00 45.16 89.37±1.71 49.42±6.66 47.84±5.22 64.63±6.17 63.96±11.29
-Smoke & Air Quality 72.46 75.38 45.48 65.96 84.20±6.11 57.79±3.33 71.38±1.20 74.93±1.17 76.24±2.51
-Support 27.27 56.52 05.86 43.84 77.72±2.26 38.76±7.23 59.99±7.05 45.73±6.62 44.34±9.21
-Other 05.22 57.58 25.36 52.46 81.29±1.00 40.44±4.48 53.68±5.21 50.58±2.45 56.33±2.27
-Weighted Average 44.69 61.17 19.86 55.94 84.48±0.69 49.30±2.01 59.56±1.10 60.90±0.80 62.65±1.29
+Tables 4 and 5 show the unimodal text/image models experimented with.
+All experiments were conducted using the same train/test split.Experiments were done on different final feature representations by using either mean pooling of the last hidden state or the CLS token.
+Freezing the first half of the encoder layers was also tried as this has been shown to increase accuracy ingle2022.
+All encoders had a linear layer appended for classification.
+For text, "roberta-base" performed best when using just the CLS token for classification and freezing the first half of the layers, as shown in Table 4.
+Later experimentation found that the unfrozen version performed better in the final classifier.
+For images, "vit-base-patch16-384" performed best when using just the CLS token for classification without any freezing.
+Many of the fusion experiments tested, as well as some architectures from other studies can be seen in Table 6.
+The same train/test split was used as before for unimodal classification.
+The best result, belonging to the selected classifier, is shown in bold.
+Table 7 shows the calculated f-scores for the VLMs, selected and baseline classifiers per class and as an average over all classes.
+Appendix I shows the confusion matrices of these classifiers.
+All metrics reported in these tables use ’weighted’ F1-score to represent the class imbalance observed in the dataset.
+Table 6: Multimodal classification results.
+Text Model Image Model Fusion Type Notes F1
+RoBERTa ViT Cross Attention encoders pretrained individually, q=image k=text v=text 0.7855
+RoBERTa ViT Cross Attention encoders pretrained individually, q=text k=image v=image 0.6489
+RoBERTa Coca Concatenative COCA/RoBERTa individually aligned w/ description of the label 0.1587
+RoBERTa+BiLSTM ViT BiLSTM+BiLinear Koshy’s model Koshy2023 0.6149
+CLIP CLIP "openai/clip-vit-base-patch32" 0.7072
+VILT VILT "dandelin/vilt-b32-finetuned-nlvr2" 0.7638
+RoBERTa ViT Concatenative ViT/RoBERTa individually aligned w/ description of the label 0.6507
+RoBERTa ViT Transformer transformer fusion: h_dim 1024 and 0.1 dropout 0.8263
+RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout 0.8364
+RoBERTa ViT Transformer RoBERTa/ViT fully frozen 0.6164
+RoBERTa ViT Concatenative stopwords removed from text before RoBERTa 0.7870
+RoBERTa ViT Concatenative CLS into single classification layer 0.8249
+RoBERTa x2 ViT Concatenative extract text from images with OCR 0.8102
+ModernBERT ViT Transformer transformer fusion: hidden dim 2048 and 0.2 dropout 0.8149
+ModernBERT ViT Concatenative CLS into single classification layer 0.7966
+RoBERTa ViT Concatenative mean-pooling of last hidden states into linear classification layer 0.7911
+RoBERTa ViT Concatenative roberta:mean, vit:cls 0.8028
+RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout RoBERTa:mean, ViT:cls 0.8223
+RoBERTa ViT Cross Attention same as proposed but with cross attention instead of transformer 0.8208
+Table 7: F1-scores per-class for zero-shot VLMs, baseline and selected classifiers.
+Baseline Classifiers Classifier zero-shot VLMs
+Class DT GNB KNN SVM w/ PCA smolVLM LLaVA Qwen GPT-4o-mini
+Evacuees 39.34±7.08 45.34±2.26 45.78±4.64 31.65±4.69 88.55±3.20 25.84 17.65 12.70 37.21
+General Information 27.04±6.43 39.11±3.73 36.01±0.56 33.64±6.32 82.04±3.54 00.00 19.05 21.88 20.37
+Preparedness 50.27±4.05 65.04±4.15 56.25±9.15 52.24±10.58 66.02±4.10 22.58 57.14 67.69 62.50
+Weather Reports 51.48±6.89 61.22±3.86 63.70±3.44 65.37±2.89 89.01±4.51 00.00 64.52 50.00 62.39
+Warnings & Status Updates 43.70±3.16 46.43±3.48 58.67±0.92 56.20±1.21 93.19±2.02 00.00 34.94 42.73 48.00
+Reports of Actions of Responders 47.50±4.05 51.49±2.81 54.84±5.75 64.06±2.15 81.10±2.14 00.00 26.09 63.89 69.09
+Infrastructure 55.13±8.21 67.51±6.28 69.16±6.58 66.01±8.15 87.88±1.86 00.00 36.36 50.00 37.50
+Political 50.04±4.69 69.18±1.74 54.64±1.51 58.89±3.39 83.73±0.29 00.00 69.09 85.25 77.86
+Insurance 86.66±1.52 91.25±1.31 90.57±3.62 92.09±1.12 87.01±1.22 93.75 88.52 90.91 86.15
+Advertisement 48.80±6.41 49.56±7.07 65.92±6.43 66.11±3.23 89.37±1.71 00.00 08.33 45.16 73.17
+Smoke and Air Quality 58.47±2.82 72.05±3.92 75.19±1.94 76.84±1.21 84.20±6.11 45.48 72.46 65.96 75.38
+Support 47.99±10.26 60.76±4.66 47.53±6.10 43.86±6.57 77.72±2.26 05.86 27.27 43.84 56.52
+Other 46.37±2.53 56.17±1.76 48.85±2.65 58.93±1.65 81.29±1.00 25.36 05.22 52.46 57.58
+F1 Weighted Average 50.88±1.11 60.77±1.45 61.22±1.23 62.21±0.44 84.48±0.69 19.86 44.69 55.94 61.17
 The custom classifier achieved the highest weighted average F1-score (84.48%), outperforming both zero-shot vision-language models and traditional baseline classifiers across all classes.
-From the results in Table 3 and Appendix J, the ‘Evacuees’ class had the lowest F1-score under the custom classifier (78.55%), likely due to high variability in both imagery and associated text.
-On the other hand, the ‘Insurance’ class achieved the highest performance (91.86%), reflecting the strong visual and textual consistency across samples.
-Other high-performing classes include ‘Political’ (89.50 ‘Smoke& Air Quality’ (89.35%), and ‘Advertisement’ (89.17%), which may be more visually or textually distinct.
-In contrast, mid-range scores are seen in classes like ‘Support’ (77.12%) and ‘Warnings & Status Updates’ (79.31%), which often contain dense textual overlays that may not align well with the image modality.
+From the results in Table 7 and Appendix I, the ‘Preparedness’ class had the lowest F1-score under the custom classifier (66.02±4.10%), likely due to high variability in both imagery and associated text.
+On the other hand, the ‘Warnings & Status Updates’ class achieved the highest performance (93.19±2.02%), reflecting the strong visual and textual consistency across samples.
+Other high-performing classes include ‘Advertisement’ (89.37±1.71) ‘Weather Reports’ (89.01±4.51%), and ‘Evacuees’ (88.55±3.20%), which may be more visually or textually distinct.
+In contrast, mid-range scores are seen in classes like ‘Support’ (77.72±2.26%) and ‘Political’ (83.73±0.29%), which often contain dense textual overlays that may not align well with the image modality.
 6 Discussion
 Task-specific fine-tuning is critical: Table 7 shows that classical models fine-tuned on labeled data generally outperform zero-shot VLMs, with the custom model achieving  23% higher f-score than VLMs.
-Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix G.
-VLMs involve tradeoffs: Zero-shot VLMs, while being a practical option when no labeled data is available, pose challenges in accuracy and cost.GPT-4o-mini offers strong results, but is costly and lacks transparency.
+Despite a modest dataset, results affirm the value of targeted annotation and model tuning, echoing findings in literature presented in Appendix F.
+VLMs involve tradeoffs: Zero-shot VLMs, while being a practical option when no labeled data is available, pose challenges in accuracy and cost.
+GPT-4o-mini offers strong results, but is costly and lacks transparency.
 Open-source models like LLaVA, Qwen, and smolVLM are more accessible but generally underperform, though Qwen is a notable exception, and trails GPT-4 by just 5%.
@@
 Future Work: The dataset exhibits a significant class imbalance, and future work may involve steps to address this issue.
-Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research Sirbu et al.
-(2022); Sîrbu et al.
-(2025).
-Fine-tuning of VLMs is also a promising avenue which should be explored in future works Duan et al.
-(2024); Zhai et al.
-(2024).
+Utilizing the unlabeled data paired with a labelled dataset in a semi-supervised setup is a promising avenue for future research sirbu2022; sirbu2025.
+Fine-tuning of VLMs is also a promising avenue which should be explored in future works duan2024cityllava; zhai2024fine.
 7 Analysis on Trends in Data
+This section describes the continuation of the analysis from the work by sherritt2025cai.
 Here, we present an analysis of an additional 46,279 posts from X, posted between 2018 and 2024, using the custom classifier described in Section 4.
@@
 This was combined with a keyword-based search, and then the data was combined, and duplicates were removed.
-The full set of search queries can be seen in Appendix K.
+The full set of search queries can be seen in Appendix J.
 Figure 2: Distribution of collected tweets based on author location.
@@
 The distribution of posts by province is shown in Figure 2, indicating that the majority of the data originates from British Columbia and Alberta.
-Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix L.
+Additional information about this data, including a word cloud of the text, the frequency distribution of the posts by word count, can be found in Appendix K.
 Using the custom classifier, predictions were obtained for the new data.
 This automated classification approach allowed for efficiency in processing the large volume of newly collected data without having to manually label each individual sample.
-This of course, will not uphold the same gold standard of labeling that a human annotator could achieve.However, while individual classifications may not always be accurate, general trends in the data can be analyzed and linked to real events.
+This of course, will not uphold the same gold standard of labeling that a human annotator could achieve.
+However, while individual classifications may not always be accurate, general trends in the data can be analyzed and linked to real events.
 Figure 3 (left) shows the total number of posts per week sorted by year.
-This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country Reuters Staff (2023); CBC News (2024).
+This demonstrates a significant event occurred the week of June 3, 2023, and July 22, 2024, due to several major fires in the country reuters2023bcbfires; cbc2024jasperfire.
 Figure 3 (right) illustrates weekly trends in predicted labels during 2023.
@@
 Warnings & Status Updates and Evacuees follow similar patterns around the week of Aug 12, which corresponds to the state of emergency declared as the city of West Kelowna was evacuated.
-Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 Williams (2023).
+Overall, at least 35,000 people were under evacuation order and another 30,000 under evacuation alert, as of August 19 williams2023bcwildfires.
 Reports of Actions of Responders and Support also rise during these peaks, reflecting coordinated emergency efforts.
@@
 Deploying these models in real-world disaster contexts without appropriate oversight, interpretability mechanisms, or human-in-the-loop verification may lead to misclassification, misinformation amplification, or missed signals, potentially affecting emergency coordination and public trust.
-Appendix A Data Collection
-Table A.1 shows the full queries used to collect the WildFireCan-MMD data from X.
-Year Query
-2022/2023 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
-2024 (#BCwildfire OR #BCfire OR #ABWildfire OR #albertawildfire OR #ABFire OR #JasperStrong OR #JasperWildfire OR #JasperAB) -has:videos has:images lang:en -is:retweet -is:quote -is:reply
-Table A.1: The queries used for collecting the data for WildFireCan-MMD.
-Appendix B Annotation
-Figure B.1 shows the distribution of the labels over the classes by each independent annotator.
-Figure B.1: Counts of labels by each annotator
-Appendix C Topic Modeling
-Figure C.1 shows examples of topics extracted by the unsupervised BERTopic.
+Appendix A Annotation
+Figure A.1 shows the distribution of the labels over the classes by each independent annotator.
+Figure A.1: Counts of labels by each annotator
+Appendix B Topic Modeling
+Figure B.1 shows examples of topics extracted by the unsupervised BERTopic.
 We present representative images and corresponding keywords generated by BERTopic for each example.
@@
 (c) Non-useful
-Figure C.1: Examples of useful/non-useful topics from topic modeling on our dataset.
-Appendix D Prompts for Zero-Shot Classification
-The full prompt used to evaluate the aforementioned VLMs is shown in Table D.1.
-Table D.1: System and User prompt used for VLMs.
+Figure B.1: Examples of useful/non-useful topics from topic modeling on our dataset.
+Appendix C Prompts for Zero-Shot Classification
+The full prompt used to evaluate the aforementioned VLMs is shown in Table C.1.
+Table C.1: System and User prompt used for VLMs.
 System Prompt:
@@
 You may only answer with the chosen option’s letter."
-Appendix E Baseline Classifiers
+Appendix D Baseline Classifiers
 To establish a baseline, several classical methods were evaluated on the dataset, namely DT, GNB, KNN, SVM.
@@
 All other hyperparameters remained default from Scikit-learn [^6] https://scikit-learn.org/stable/.
-Five-fold cross validation was used to train and test all methods to report a confidence interval.
-Appendix F VLMs as Zero-Shot Classifiers:
-Four VLMs were considered for experimentation, GPT-4o-mini Achiam et al.
-(2023), LLaVA Liu et al.
-(2023), smolVLM Marafioti et al.
-(2025), Qwen2.5-VL Bai et al.
-(2025) and tested them in a zero-shot setting.
-Classification prompts for all three VLMs were created.
+Three different random splits were used to train and test all methods to report a confidence interval.
+Appendix E VLMs as Zero-Shot Classifiers:
+Four VLMs were considered for experimentation, GPT-4o-mini OpenAI2023, LLaVA Liu2023, smolVLM smolvlm2025, Qwen2.5-VL qwen2025 and tested in a zero-shot setting.
+Classification prompts for all four VLMs were created.
 With each query to the model, a list of all categories and their descriptions are given as a prompt along with a data instance (text + image) to classify.
 The prompt specifies that the model’s task is to choose one of the provided categories and respond only with its corresponding letter.
-The system and user prompts can be seen in Appendix D and are the same for all VLMs.
+The system and user prompts can be seen in Appendix C and are the same for all VLMs.
 The same hyperparameters are used for all VLMs with temperature=0.1 num_beams=1, max_new_tokens=1024.
-The same test set was used for all VLMs and consisted of a stratified split with random_state=8.
-This single split was discovered to be a good split from previous experimentation with baseline and custom classifiers, and was necessary due to the cost of running the VLM experiments multiple times.
+The same test set was used for all VLMs and consisted of a stratified split.
+This single split was discovered to be a good split from previous experimentation with baseline and custom classifiers, and is one of the same three splits used for these classifiers.
 SmolVLM-2B: Being the smallest of the VLMs experimented with, with only 2.2B parameters and <10GB of GPU ram, experiments were able to be performed locally without quantization.
@@
 Being too large for the hardware available, manual quantization using BitsandBytes was performed to shrink the precision down to 4-bit to be used locally.
-LLaVA-13B: Also considered an intermediately-sized option, the 13B 4-bit quantized version of LLaVA was used due to hardware constraints.
+LLaVA-v1.5-13B: Also considered an intermediately-sized option, the 13B 4-bit quantized version of LLaVA was used due to hardware constraints.
 Using a combination of a publicly available Colab notebook  [^7] https://github.com/camenduru/LLaVA-colab, the author of which performed the quantization, and the original code from the open-source LLaVA repository  [^8] https://github.com/haotian-liu/LLaVA, the model was able to be used with Google Colab.
 GPT-4o-mini: The largest of the VLMs tested, a paid API service had to be used for querying GPT-4o-mini 999version:GPT-4o-mini-2024-07-18, and it is also the only closed-source model tested.
-A specified content field in each query contains the system prompt, text and image from the tweet, and user prompt.
-For each method, the responses are saved as a file where each line contains the model’s answer supplied as a letter from the options shown in Table D.1.
+For each method, the responses are saved as a file where each line contains the model’s answer supplied as a letter from the options shown in Table C.1.
 This then allows for the calculation of metrics.
-Appendix G Related Architectures
-Table G.1 briefly describes the architectures proposed by various authors on the CrisisMMD dataset.
+Appendix F Related Architectures
+Table F.1 briefly describes the architectures proposed by various authors for the CrisisMMD dataset.
 The purpose of this table is to demonstrate the literature review used to select the classifier proposed in section 4 and give an idea of the expected accuracy of such a classifier on this task of wildfire social-media data.
 Paper Method Task 1 Task 2 Task 3
-Agar. 2020 Agarwal et al. (2020) InceptionV3+RCNN+AttentionFusion 99.00 99.00 97.00
-Abav. 2020 Abavisani et al. (2020) DenseNet+BERT+CrossAttentionFusion 89.35 91.82 70.41
-Ofli. 2020 Ofli et al. (2020) VGG16+CNN+Concat.Fusion 84.20 78.30 -
-Xukun 2020 Li and Caragea (2020) AlexNet+GRU+Concat.Fusion 78.13 - -
-Madi. 2021 Madichetty et al. (2021) DenseNet+BERT+AverageFusion 74.85 - -
-Kumar 2022 Kumar et al. (2022) VGG16+LSTM+Concat.Fusion 82.00 - -
-Pran. 2022 Pranesh (2022) ResNet+RoBERTa+TransformerAttentionFusion 85.10 89.50 -
-Sirbu 2022 Sirbu et al. (2022) semi-supervised 82.00 87.20 88.70
-Diva. 2022 Divakaran et al. (2022) DenseNet+RoBERTa+Concat.Fusion 89.26 85.87 -
-Kotha 2022 Kotha et al. (2022) RegNetY320+BERT+Concat.Fusion 89.63 86.59 -
-Khat. 2022 Khattar and Quadri (2022) VGG16+BiLSTM+CrossAttentionFusion 84.08 - -
-Madi. 2023 Madichetty et al. (2023) VGG16+RoBERTa+MultiplicativeFusion 90.21 - -
-Basit 2023 Basit et al. (2023) DeIT+ALBERT+Concat.Fusion 79.00 80.00 86.00
-Koshy 2023 Koshy and Elango (2023) ViT+RoBERTa+BiLSTM+BilinearFusion 97.14 - -
-Gupta 2024 Gupta et al. (2024) DenseNet+ELECTRA+CrossAttentionFusion 91.20 93.40 72.20
-Shet. 2025 Shetty et al. (2025) ViT+RoBERTa+CrossAttentionFusion 91.25 89.63 68.72
-Sirbu 2025 Sîrbu et al. (2025) semi-supervised 91.22 89.55 -
-Teng 2025 Teng and Öhman (2025) VGG16+BERT+CrossAttentionFusion 81.00 - -
-Dar 2025 Dar et al. (2025) (ResNet50+BERT)+(Graph)+IDEAFusion 98.23 90.13 -
-Table G.1: Literature review of deep-learning approaches used by other works on the CrisisMMD dataset.
-Appendix H Ablation & Other Experiments
-Tables H.1 and H.2 show the different text/image encoders experimented with, using either only text or only images respectively.
-All experiments here were conducted using the same train/test split, which was earlier identified as being a good split through cross validation experiments on different random seeds.
-Experiments were done on different final feature representations by using either mean pooling of the last hidden state or, using the CLS token.
-Also, freezing the first half of the layers as this has been shown to increase accuracy Ingle et al.
-(2022).
-All encoders had a linear layer appended for classification.
-For text, "roberta-base" performed best when using just the CLS token for classification and freezing the first half of the encoder layers, as shown in Table H.1.
-Later experimentation found that the unfrozen version performed better in the final classifier.
-For images, "vit-base-patch16-384" performed best when using just the CLS token for classification without any freezing.
-Many of the fusion experiments tested, as well as some architectures from other studies can be seen in Table H.3.
-The same train/test split was used as before for unimodal classification.
-The best result, belonging to the selected classifier, is shown in bold.
-Table H.4 shows the ablation study performed on the selected model.
-(1) is the result of the selected model, (2) is the result of the model with a simpler classification head consisting of a single linear layer, (3) is the model with the proposed classification head but without the fusion module,
-Three-fold cross-validation was performed later to support the result as described in section 4.
-Text Model Notes F1-Weighted
-’bert-base-cased’ mean-pooling 0.7779
-’bert-base-cased’ CLS 0.7793
-"bertweet-base" mean-pooling 0.7900
-"bertweet-base" CLS 0.7927
-"ModernBERT-base" mean-pooling 0.7934
-"ModernBERT-base" CLS 0.7862
-"roberta-base" robertaforseq 0.7718
-"roberta-base" mean-pooling 0.7934
-"roberta-base" mean-pooling, half-frozen 0.7940
-"roberta-base" mean-pooling, custom vocab 0.7209
-"roberta-base" CLS 0.7798
-"roberta-base" CLS, half-frozen 0.8008
-Table H.1: Text-only results.
-Image Model Notes F1-Weighted
-DenseNet121 ImageNet weights 0.5733
-InceptionV3 ImageNet weights 0.5804
-ResNet50 ImageNet weights 0.5841
-VGG16 ImageNet weights 0.5993
-"facebook/deit-base-patch16-384" CLS 0.6260
-"google/vit-base-patch16-384" half-frozen, CLS 0.6290
-"google/vit-base-patch16-384" mean-pooling 0.6404
-"google/vit-base-patch16-384" CLS 0.6419
-Table H.2: Image-only results.
-Text Model Image Model Fusion Type Notes F1
-RoBERTa ViT Cross Attention encoders pretrained individually, q=image k=text v=text 0.7855
-RoBERTa ViT Cross Attention encoders pretrained individually, q=text k=image v=image 0.6489
-RoBERTa Coca Concatenative COCA/RoBERTa individually aligned w/ description of the label 0.1587
-RoBERTa+BiLSTM ViT BiLSTM+BiLinear Koshy’s model Koshy and Elango (2023) 0.6149
-CLIP CLIP "openai/clip-vit-base-patch32" 0.7072
-VILT VILT "dandelin/vilt-b32-finetuned-nlvr2" 0.7638
-RoBERTa ViT Concatenative ViT/RoBERTa individually aligned w/ description of the label 0.6507
-RoBERTa ViT Transformer transformer fusion: h_dim 1024 and 0.1 dropout 0.8263
-RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout 0.8364
-RoBERTa ViT Transformer RoBERTa/ViT fully frozen 0.6164
-RoBERTa ViT Concatenative stopwords removed from text before RoBERTa 0.7870
-RoBERTa ViT Concatenative CLS into single classification layer 0.8249
-RoBERTa x2 ViT Concatenative extract text from images with OCR 0.8102
-ModernBERT ViT Transformer transformer fusion: hidden dim 2048 and 0.2 dropout 0.8149
-ModernBERT ViT Concatenative CLS into single classification layer 0.7966
-RoBERTa ViT Concatenative mean-pooling of last hidden states into linear classification layer 0.7911
-RoBERTa ViT Concatenative roberta:mean, vit:cls 0.8028
-RoBERTa ViT Transformer transformer fusion: h_dim 2048 and 0.2 dropout RoBERTa:mean, ViT:cls 0.8223
-RoBERTa ViT Cross Attention same as proposed but with cross attention instead of transformer 0.8208
-Table H.3: Multimodal results.
+Agar. 2020 Agarwal2020 InceptionV3+RCNN+AttentionFusion 99.00 99.00 97.00
+Abav. 2020 Abavisani2020 DenseNet+BERT+CrossAttentionFusion 89.35 91.82 70.41
+Ofli. 2020 Ofli2020 VGG16+CNN+Concat.Fusion 84.20 78.30 -
+Xukun 2020 xukun2020 AlexNet+GRU+Concat.Fusion 78.13 - -
+Madi. 2021 Madichetty2021 DenseNet+BERT+AverageFusion 74.85 - -
+Kumar 2022 kumar2022 VGG16+LSTM+Concat.Fusion 82.00 - -
+Pran. 2022 pranesh2022 ResNet+RoBERTa+TransformerAttentionFusion 85.10 89.50 -
+Sirbu 2022 sirbu2022 semi-supervised 82.00 87.20 88.70
+Diva. 2022 Divakaran2022 DenseNet+RoBERTa+Concat.Fusion 89.26 85.87 -
+Kotha 2022 kotha2022 RegNetY320+BERT+Concat.Fusion 89.63 86.59 -
+Khat. 2022 Khattar2022 VGG16+BiLSTM+CrossAttentionFusion 84.08 - -
+Madi. 2023 Madichetty2023 VGG16+RoBERTa+MultiplicativeFusion 90.21 - -
+Basit 2023 Basit2023 DeIT+ALBERT+Concat.Fusion 79.00 80.00 86.00
+Koshy 2023 Koshy2023 ViT+RoBERTa+BiLSTM+BilinearFusion 97.14 - -
+Gupta 2024 Gupta2024 DenseNet+ELECTRA+CrossAttentionFusion 91.20 93.40 72.20
+Shet. 2025 shetty2025 ViT+RoBERTa+CrossAttentionFusion 91.25 89.63 68.72
+Sirbu 2025 sirbu2025 semi-supervised 91.22 89.55 -
+Teng 2025 teng2025 VGG16+BERT+CrossAttentionFusion 81.00 - -
+Dar 2025 Dar2025 (ResNet50+BERT)+(Graph)+IDEAFusion 98.23 90.13 -
+Table F.1: Literature review of deep-learning approaches used by other works on the CrisisMMD dataset.
+Appendix G Ablation & Other Experiments
+Table G.1 shows the ablation study performed on the selected model.
+(1) is the result of the selected model, (2) is the result of the model with a simpler classification head consisting of a single linear layer, (3) is the model with the proposed classification head but without the fusion module.
+Two other random splits were later used to support the result described in section 4.
 Ablation Explanation F1-Weighted
@@
 3 No fusion module 81.71
-Table H.4: Ablation of model components.
-Appendix I Structure of Custom Classifier
-Figure I.1 shows the architecture of the selected approach. This model is described in detail in Section 4.
-Figure I.1: Selected dual-encoder architecture utilizing transformer fusion module.
-Appendix J Confusion Matrices
-For deeper insights into the classifiers performances, the confusion matrices are shown in Figures J.1 and J.2.
+Table G.1: Ablation of model components.
+Appendix H Structure of Custom Classifier
+Figure H.1 shows the architecture of the selected approach. This model is described in detail in Section 4.
+Figure H.1: Selected dual-encoder architecture utilizing transformer fusion module.
+Appendix I Confusion Matrices
+For deeper insights into the classifiers performances, the confusion matrices are shown in Figures I.1 and I.2.
 (a) Random seed: 8
@@
 (c) Random seed: 14
-Figure J.1: Confusion matricies for selected classifier over the different train/test splits.
+Figure I.1: Confusion matricies for selected classifier over the different train/test splits.
 (a) LLaVA.
@@
 (d) Qwen.
-Figure J.2: Confusion matrices of VLMs.
-Appendix K Queries Used for Additional Data Collection
+Figure I.2: Confusion matrices of VLMs.
+Appendix J Queries Used for Additional Data Collection
 Shown here is the list of queries and keywords used to collect the additional unlabeled data discussed in Section 7.
@@
 Note: Every query includes the suffix -has:videos has:images lang:en -is:retweet -is:quote -is:reply to filter for original English-language posts with images only.
-Appendix L Unlabeled Dataset
+Appendix K Unlabeled Dataset
 Figure 2 shows the distribution of post origins.
-A word cloud of the text data of the unlabeled dataset collected in Section 7, and also the frequency distribution of the posts by word count, are shown in Figure L.1.
+A word cloud of the text data of the unlabeled dataset collected in Section 7, and also the frequency distribution of the posts by word count, are shown in Figure K.1.
 (a) Word cloud of the posts.
 (b) Frquency distribution of the posts w.r.t their lengths.
-Figure L.1: (a) Word cloud of the text from the unlabeled tweets and (b) the frequency distribution of their word counts.
+Figure K.1: (a) Word cloud of the text from the unlabeled tweets and (b) the frequency distribution of their word counts.